### v3

# Is Chain-of-Thought Reasoning of LLMs a Mirage? A Data Distribution Lens

Abstract
Chain-of-Thought (CoT) prompting has been shown to improve Large Language Model (LLM) performance on various tasks. With this approach, LLMs appear to produce human-like reasoning steps before providing answers (a.k.a., CoT reasoning), which often leads to the perception that they engage in deliberate inferential processes. However, some initial findings suggest that CoT reasoning may be more superficial than it appears, motivating us to explore further. In this paper, we study CoT reasoning via a data distribution lens and investigate if CoT reasoning reflects a structured inductive bias learned from in-distribution data, allowing the model to conditionally generate reasoning paths that approximate those seen during training. Thus, its effectiveness is fundamentally bounded by the degree of distribution discrepancy between the training data and the test queries. With this lens, we dissect CoT reasoning via three dimensions: task, length, and format. To investigate each dimension, we design DataAlchemy, an isolated and controlled environment to train LLMs from scratch and systematically probe them under various distribution conditions. Our results reveal that CoT reasoning is a brittle mirage that vanishes when it is pushed beyond training distributions. This work offers a deeper understanding of why and when CoT reasoning fails, emphasizing the ongoing challenge of achieving genuine and generalizable reasoning. Our code is available at GitHub: https://github.com/ChengshuaiZhao0/DataAlchemy.
1 Introduction
Figure 1: The data perspective lens. CoT reasoning’s effectiveness is fundamentally bounded by the degree of distribution discrepancy between the training data and the test queries. Guided by this lens, we dissect CoT reasoning via three dimensions: task, length, and format.
Recent years have witnessed Large Language Models’ (LLMs) dominant role in various domains (Zhao et al., 2023; Li et al., 2025b; Zhao et al., 2025; Ting et al., 2025) through versatile prompting techniques (Wei et al., 2022; Yao et al., 2023; Kojima et al., 2022). Among these, Chain-of-Thought (CoT) prompting (Wei et al., 2022) has emerged as a prominent method for eliciting structured reasoning from LLMs (a.k.a., CoT reasoning). By appending a simple cue such as “Let’s think step by step,” LLMs decompose complex problems into intermediate steps, producing outputs that resemble human-like reasoning. It has been shown to be effective in tasks requiring logical inference(Xu et al., 2024), mathematical problem solving (Imani et al., 2023), and commonsense reasoning (Wei et al., 2022). The empirical successes of CoT reasoning lead to the perception that LLMs engage in deliberate inferential processes (Yu et al., 2023; Zhang et al., 2024a; Ling et al., 2023; Zhang et al., 2024c).
However, a closer examination reveals inconsistencies that challenge this optimistic view. Consider this straightforward question: “The day the US was established is in a leap year or a normal year?” When prompted with the CoT prefix, the modern LLM Gemini responded: “The United States was established in 1776. 1776 is divisible by 4, but it’s not a century year, so it’s a leap year. Therefore, the day the US was established was in a normal year.” This response exemplifies a concerning pattern: the model correctly recites the leap year rule and articulates intermediate reasoning steps, yet produces a logically inconsistent conclusion (i.e., asserting 1776 is both a leap year and a normal year). Such inconsistencies suggest that there is a distinction between human-like inference and CoT reasoning.
An expanding body of analyses reveals that LLMs tend to rely on surface-level semantics and clues rather than logical procedures (Stechly et al., 2024; Kambhampati, 2024; Chen et al., 2025b; Lanham et al., 2023). LLMs construct superficial chains of logic based on learned token associations, often failing on tasks that deviate from commonsense heuristics or familiar templates (Tang et al., 2023). In the reasoning process, performance degrades sharply when irrelevant clauses are introduced, which indicates that models cannot grasp the underlying logic (Mirzadeh et al., 2024). This fragility becomes even more apparent when models are tested on more complex tasks, where they frequently produce incoherent solutions and fail to follow consistent reasoning paths (Shojaee et al., 2025). Collectively, these pioneering works deepen the skepticism surrounding the true nature of CoT reasoning.
In light of this line of research, we question the CoT reasoning by proposing an alternative lens through data distribution and further investigating why and when it fails. We hypothesize that CoT reasoning reflects a structured inductive bias learned from in-distribution data, allowing the model to conditionally generate reasoning paths that approximate those seen during training. As such, its effectiveness is inherently limited by the nature and extent of the distribution discrepancy between training data and the test queries. Guided by this data distribution lens, we dissect CoT reasoning via three dimensions: (i) task—To what extent CoT reasoning can handle tasks that involve transformations or previously unseen task structures. (2) length—how CoT reasoning generalizes to chains with length different from that of training data; and (3) format—how sensitive CoT reasoning is to surface-level query form variations. To evaluate each aspect, we introduce DataAlchemy, a controlled and isolated experiment that allows us to train LLMs from scratch and systematically probe them under various distribution shifts.
Our findings reveal that CoT reasoning works effectively when applied to in-distribution or near in-distribution data but becomes fragile and prone to failure even under moderate distribution shifts. In some cases, LLMs generate fluent yet logically inconsistent reasoning steps. The results suggest that what appears to be structured reasoning can be a mirage, emerging from memorized or interpolated patterns in the training data rather than logical inference. These insights carry important implications for both practitioners and researchers. For practitioners, our results highlight the risk of relying on CoT as a plug-and-play solution for reasoning tasks and caution against equating CoT-style output with human thinking. For researchers, the results underscore the ongoing challenge of achieving reasoning that is both faithful and generalizable, motivating the need to develop models that can move beyond surface-level pattern recognition to exhibit deeper inferential competence. Our contributions are summarized as follows:
  • $\star$
    Novel perspective. We propose a data distribution lens for CoT reasoning, illuminating that its effectiveness stems from structured inductive biases learned from in-distribution training data. This framework provides a principled lens for understanding why and when CoT reasoning succeeds or fails.
  • $\star$
    Controlled environment. We introduce DataAlchemy, an isolated experimental framework that enables training LLMs from scratch and systematically probing CoT reasoning. This controlled setting allows us to isolate and analyze the effects of distribution shifts on CoT reasoning without interference from complex patterns learned during large-scale pre-training.
  • $\star$
    Empirical validation. We conduct systematic empirical validation across three critical dimensions—task, length, and format. Our experiments demonstrate that CoT reasoning exhibits sharp performance degradation under distribution shifts, revealing that seemingly coherent reasoning masks shallow pattern replication.
  • $\star$
    Real-world implication. This work reframes the understanding of contemporary LLMs’ reasoning capabilities and emphasizes the risk of over-reliance on COT reasoning as a universal problem-solving paradigm. It underscores the necessity for proper evaluation methods and the development of LLMs that possess authentic and generalizable reasoning capabilities.
2 Related Work
2.1 LLM Prompting and CoT
Chain-of-Thought (CoT) prompting revolutionized how we elicit reasoning from Large Language Models by decomposing complex problems into intermediate steps (Wei et al., 2022). By augmenting few-shot exemplars with reasoning chains, CoT showed substantial performance gains on various tasks (Xu et al., 2024; Imani et al., 2023; Wei et al., 2022). Building on this, several variants emerged. Zero-shot CoT triggers reasoning without exemplars using instructional prompts (Kojima et al., 2022), and self-consistency enhances performance via majority voting over sampled chains (Wang et al., 2023). To reduce manual effort, Auto-CoT generates CoT exemplars using the models themselves (Zhang et al., 2023). Beyond linear chains, Tree-of-Thought (ToT) frames CoT as a tree search over partial reasoning paths (Yao et al., 2023), enabling lookahead and backtracking. SymbCoT combines symbolic reasoning with CoT by converting problems into formal representations (Xu et al., 2024). Recent work increasingly integrates CoT into the LLM inference process, generating long-form CoTs (Jaech et al., 2024; Team, 2024; Guo et al., 2025; Team et al., 2025). This enables flexible strategies like mistake correction, step decomposition, reflection, and alternative reasoning paths (Yeo et al., 2025; Chen et al., 2025a). The success of prompting techniques and long-form CoTs has led many to view them as evidence of emergent, human-like reasoning in LLMs. In this work, we challenge that viewpoint by adopting a data-centric perspective and demonstrating that CoT behavior arises largely from pattern matching over training distributions.
2.2 Discussion on Illusion of LLM Reasoning
While Chain-of-Thought prompting has led to impressive gains on complex reasoning tasks, a growing body of work has started questioning the nature of these gains (Stechly et al., 2024, 2025; Kambhampati et al., 2025). One major line of research highlights the fragility of CoT reasoning. Minor and semantically irrelevant perturbations such as distractor phrases or altered symbolic forms can cause significant performance drops in state-of-the-art models (Mirzadeh et al., 2024; Tang et al., 2023). Models often incorporate such irrelevant details into their reasoning, revealing a lack of sensitivity to salient information. Other studies show that models prioritize the surface form of reasoning over logical soundness; in some cases, longer but flawed reasoning paths yield better final answers than shorter, correct ones (Bentham et al., 2024). Similarly, performance does not scale with problem complexity as expected—models may overthink easy problems and give up on harder ones (Shojaee et al., 2025). Another critical concern is the faithfulness of the reasoning process. Intervention-based studies reveal that final answers often remain unchanged even when intermediate steps are falsified or omitted (Lanham et al., 2023), a phenomenon dubbed the illusion of transparency (Bentham et al., 2024; Chen et al., 2025b). Together, these findings suggest that LLMs are not principled reasoners but rather sophisticated simulators of reasoning-like text. However, a systematic understanding of why and when CoT reasoning fails is still a mystery.
2.3 OOD Generalization of LLMs
Out-of-distribution (OOD) generalization, where test inputs differ from training data, remains a key challenge in machine learning, particularly for large language models (LLMs)(Yang et al., 2024, 2023; Budnikov et al., 2025; Zhang et al., 2024b). Recent studies show that LLMs prompted to learn novel functions often revert to similar functions encountered during pretraining (Wang et al., 2024; Garg et al., 2022). Likewise, LLM generalization frequently depends on mapping new problems onto familiar compositional structures (Song et al., 2025). CoT prompting improves OOD generalization (Wei et al., 2022), with early work demonstrating length generalization for multi-step problems beyond training distributions (Yao et al., 2025; Shen et al., 2025). However, this ability is not inherent to CoT and heavily depends on model architecture and training setups. For instance, strong generalization in arithmetic tasks was achieved only when algorithmic structures were encoded into positional encodings (Cho et al., 2024). Similarly, finer-grained CoT demonstrations during training boost OOD performance, highlighting the importance of data granularity (Wang et al., 2025a). Theoretical and empirical evidence shows that CoT generalizes well only when test inputs share latent structures with training data; otherwise, performance declines sharply (Wang et al., 2025b; Li et al., 2025a). Despite its promise, CoT still struggles with genuinely novel tasks or formats. In the light of these brilliant findings, we propose rethinking CoT reasoning through a data distribution lens: decomposing CoT into task, length, and format generalization, and systematically investigating each in a controlled setting.
Figure 2: Framework of DataAlchemy. It creates an isolated and controlled environment to train LLMs from scratch and probe the task, length, and format generalization.
3 The Data Distribution Lens
We propose a fundamental reframing to understand what CoT actually represents. We hypothesize that the underlying mechanism is better understood through the lens of data distribution: rather than executing explicit reasoning procedures, CoT operates as a pattern-matching process that interpolates and extrapolates from the statistical regularities present in its training distribution. Specifically, we posit that CoT’s success stems not from a model’s inherent reasoning capacity, but from its ability to generalize conditionally to out-of-distribution (OOD) test cases that are structurally similar to in-distribution exemplars.
To formalize this view, we model CoT prompting as a conditional generation process constrained by the distributional properties of the training data. Let $\mathcal{D}_{\text{train}}$ denote the training distribution over input-output pairs $(x,y)$, where $x$ represents a reasoning problem and $y$ denotes the solution sequence (including intermediate reasoning steps). The model learns an approximation $f_{\theta}(x)\approx y$ by minimizing empirical risk over samples drawn from $\mathcal{D}_{\text{train}}$.
Let the expected training risk be defined as:
$R_{\text{train}}(f_{\theta})=\mathbb{E}_{(x,y)\sim\mathcal{D}_{\text{train}}}[\ell(f_{\theta}(x),y)],$ (1)
where $\ell$ is a task-specific loss function (e.g., cross-entropy, token-level accuracy). At inference time, given a test input $a_{\text{test}}$ sampled from a potentially different distribution $\mathcal{D}_{\text{test}}$, the model generates a response $y_{\text{test}}$ conditioned on patterns learned from $\mathcal{D}_{\text{train}}$. The corresponding expected test risk is:
$R_{\text{test}}(f_{\theta})=\mathbb{E}_{(x,y)\sim\mathcal{D}_{\text{test}}}[\ell(f_{\theta}(x),y)].$ (2)
The degree to which the model generalizes from $\mathcal{D}_{\text{train}}$ to $\mathcal{D}_{\text{test}}$ is governed by the distributional discrepancy between the two, which we quantify using divergence measures:
Definition 3.1 (Distributional Discrepancy).
Given training distribution $\mathcal{D}_{\text{train}}$ and test distribution $\mathcal{D}_{\text{test}}$, the distributional discrepancy is defined as:
$\Delta(\mathcal{D}_{\text{train}},\mathcal{D}_{\text{test}})=\mathcal{H}(\mathcal{D}_{\text{train}}\parallel\mathcal{D}_{\text{test}})$ (3)
where $\mathcal{H}(\cdot\parallel\cdot)$ is a divergence measure (e.g., KL divergence, Wasserstein distance) that quantifies the statistical distance between the two distributions.
Theorem 3.1 (CoT Generalization Bound).
Let $f_{\theta}$ denote a model trained on $\mathcal{D}_{\text{train}}$ with expected training risk $R_{\text{train}}(f_{\theta})$. For a test distribution $\mathcal{D}_{\text{test}}$, the expected test risk $R_{\text{test}}(f_{\theta})$ is bounded by:
$R_{\text{test}}(f_{\theta})\leq R_{\text{train}}(f_{\theta})+\Lambda\cdot\Delta(\mathcal{D}_{\text{train}},\mathcal{D}_{\text{test}})+\mathcal{O}\left(\sqrt{\frac{\log(1/\delta)}{n}}\right)$ (4)
where $\Lambda>0$ is a Lipschitz constant that depends on the model architecture and task complexity, $n$ is the training sample size, and the bound holds with probability $1-\delta$, where $\delta$ is the failure probability.
The proof is provided in Appendix A.1
Building on this data distribution perspective, we identify three critical dimensions along which distributional shifts can occur, each revealing different aspects of CoT’s pattern-matching nature: ➊ Task generalization examines how well CoT transfers across different types of reasoning tasks. Novel tasks may have unique elements and underlying logical structure, which introduces distributional shifts that challenge the model’s ability to apply learned reasoning patterns. ➋ Length generalization investigates CoT’s robustness to reasoning chains of varying lengths. Since training data typically contains reasoning sequences within a certain length range, test cases requiring substantially longer or shorter reasoning chains represent a form of distributional shift along the sequence length dimension. This length discrepancy could result from the reasoning step or the text-dependent solution space. ➌ Format generalization explores how sensitive CoT is to variations in prompt formulation and structure. Due to various reasons (e.g., sophistical training data or diverse background of users), it is challenging for LLM practitioners to design a golden prompt to elicit knowledge suitable for the current case. Their detailed definition and implementation are given in subsequent sections.
Each dimension provides a unique lens for understanding the boundaries of CoT’s effectiveness and the mechanisms underlying its apparent reasoning capabilities. By systematically varying these dimensions in controlled experimental settings, we can empirically validate our hypothesis that CoT performance degrades predictably as distributional discrepancy increases, thereby revealing its fundamental nature as a pattern-matching rather than reasoning system.
4 DataAlchemy: An Isolated and Controlled Environment
To systematically investigate the influence of distributional shifts on CoT reasoning capabilities, we introduce DataAlchemy, a synthetic dataset framework designed for controlled experimentation. This environment enables us to train language models from scratch under precisely defined conditions, allowing for rigorous analysis of CoT behavior across different OOD scenarios. The overview is shown in Figure 2.
4.1 Basic Atoms and Elements
Let $\mathcal{A}=\{\texttt{A},\texttt{B},\texttt{C},\ldots,\texttt{Z}\}$ denote the alphabet of 26 basic atoms. An element $\mathbf{e}$ is defined as an ordered sequence of atoms:
$\mathbf{e}=(a_{0},a_{1},\ldots,a_{l-1})\quad\text{where}\quad a_{i}\in\mathcal{A},\quad l\in\mathbb{Z}^{+}$ (5)
This design provides a versatile manipulation for the size of the dataset $\mathcal{D}$ (i.e., $|\mathcal{D}|=|\mathcal{A}|^{l}$) by varying element length $l$ to train language models with various capacities. Meanwhile, it also allows us to systematically probe text length generalization capabilities.
4.2 Transformations
A transformation is an operation that operates on elements $F:\mathbf{e}\rightarrow\hat{\mathbf{e}}$. In this work, we consider two fundamental transformations: the ROT Transformation and the Cyclic Position Shift. To formally define the transformations, we introduce a bijective mapping $\phi:\mathcal{A}\to\mathbb{Z}_{26}$, where $\mathbb{Z}_{26}=\{0,1,\ldots,25\}$, such that $\phi(c)$ maps a character to its zero-based alphabetical index.
Definition 4.1 (ROT Transformation).
Given an element $\mathbf{e}=(a_{0},\ldots,a_{l-1})$ and a rotation parameter $n\in\mathbb{Z}$, the ROT Transformation $f_{\text{rot}}$ produces an element $\hat{\mathbf{e}}=(\hat{a}_{0},\ldots,\hat{a}_{l-1})$. Each atom $\hat{a}_{i}$ is:
$\hat{a}_{i}=\phi^{-1}((\phi(a_{i})+n)\pmod{26})$ (6)
This operation cyclically shifts each atom $n$ positions forward in alphabetical order. For example, if $\mathbf{e}=(\texttt{A},\texttt{P},\texttt{P},\texttt{L},\texttt{E})$ and $n=13$, then $f_{\text{rot}}(\mathbf{e},13)=(\texttt{N},\texttt{C},\texttt{C},\texttt{Y},\texttt{R})$.
Definition 4.2 (Cyclic Position Shift).
Given an element $\mathbf{e}=(a_{0},\ldots,a_{l-1})$ and a shift parameter $n\in\mathbb{Z}$, the Cyclic Position Shift $f_{\text{pos}}$ produces an element $\hat{\mathbf{e}}=(\hat{a}_{0},\ldots,\hat{a}_{l-1})$. Each atom $\hat{a}_{i}$ is defined by a cyclic shift of indices:
$\hat{a}_{i}=a_{(i-n)\pmod{l}}$ (7)
This transformation cyclically shifts the positions of the atoms within the sequence by $n$ positions to the right. For instance, if $\mathbf{e}=(\texttt{A},\texttt{P},\texttt{P},\texttt{L},\texttt{E})$ and $n=1$, then $f_{\text{pos}}(\mathbf{e},1)=(\texttt{E},\texttt{A},\texttt{P},\texttt{P},\texttt{L})$.
Definition 4.3 (Generalized Compositional Transformation).
To model multi-step reasoning, we define a compositional transformation as the successive application of a sequence of operations. Let $S=(f_{1},f_{2},\ldots,f_{k})$ be a sequence of operations, where each $f_{i}$ is one of the fundamental transformations $\mathcal{F}=\{f_{\text{rot}},f_{\text{pos}}\}$ with its respective parameters. The compositional transformation $f_{\text{S}}$ for the sequence $S$ is the function composition:
$f_{\text{S}}=f_{k}\circ f_{k}\circ\cdots\circ f_{1}$ (8)
The resulting element $\hat{\mathbf{e}}$ is obtained by applying the operations sequentially to an initial element $\mathbf{e}$:
$\hat{\mathbf{e}}=f_{k}(f_{k-1}(\ldots(f_{1}(\mathbf{e}))\ldots))$ (9)
This design enables the construction of arbitrarily complex transformation chains by varying the type, parameters, order, and length of operations within the sequence. At the sample time, we can naturally acquire the COT reasoning step by decomposing the intermediate process:
$\underbrace{f_{\text{S}}(\mathbf{e}):}_{\text{Query}}\quad\underbrace{\mathbf{e}\xrightarrow{f_{1}}\mathbf{e}^{(1)}\xrightarrow{f_{2}}\mathbf{e}^{(2)}\cdots\xrightarrow{f_{k-1}}\mathbf{e}^{(k-1)}\xrightarrow{f_{k}}}_{\text{COT reasoning steps}}\underbrace{\boxed{\hat{\mathbf{e}}}}_{\text{Answer}}$ (10)
4.3 Environment Setting
Through systematic manipulation of elements and transformations, DataAlchemyoffers a flexible and controllable framework for training LLMs from scratch, facilitating rigorous investigation of diverse OOD scenarios. Without specification, we employ a decoder-only language model GPT-2 (Radford et al., 2019) with a configuration of 4 layers, 32 hidden dimensions, and 4 attention heads. We utilize a Byte-Pair Encoding (BPE) tokenizer. Both LLMs and the tokenizer follow the general modern LLM pipeline. During the inference time, we set the temperature to 1e-5. For rigor, we also study LLMs with various parameters, architectures, and temperatures in Section 8. Details of the implementation are provided in the Appendix B. We consider that each element consists of 4 basic atoms, which produces 456,976 samples for each dataset with varied transformations and token amounts. We initialize the two transformations $f_{1}=f_{\text{rot}}(e,13)$ and $f_{2}=f_{\text{pos}}(e,1)$. We consider the exact match rate, Levenshtein distance (i.e., edit distance) (Yujian and Bo, 2007), and BLEU score (Papineni et al., 2002) as metrics and evaluate the produced reasoning step, answer, and full chain. Examples of the datasets and evaluations are shown in Appendix C
5 Task Generalization
Task generalization represents a fundamental challenge for CoT reasoning, as it directly tests a model’s ability to apply learned concepts and reasoning patterns to unseen scenarios. In our controlled experiments, both transformation and elements could be novel. Following this, we decompose task generalization into two primary dimensions: element generalization and transformation generalization.
Task Generalization Complexity. Guided by the data distribution lens, we first introduce a measure for generalization difficulty:
Proposition 5.1 (Task Generalization Complexity).
For a reasoning chain $f_{S}$ operating on elements $\mathbf{e}=(a_{0},\ldots,a_{l-1})$, define:
$\operatorname{TGC}(C)=$ $\alpha\sum_{i=1}^{m}\mathbb{I}\left[a_{i}\notin\mathcal{E}^{i}_{\text{train }}\right]+\beta\sum_{j=1}^{n}\mathbb{I}\left[f_{j}\notin\mathcal{F}_{\text{train }}\right]+\gamma\mathbb{I}\left[\left(f_{1},f_{2},\ldots,f_{k}\right)\notin\mathcal{P}_{\text{train }}\right]+C_{T}$ (11)
[…]
We establish a critical threshold beyond which CoT reasoning fails exponentially:
Theorem 5.1 (Task Generalization Failure Threshold).
There exists a threshold $\tau$ such that when $\text{TGC}(C)>\tau$, the probability of correct CoT reasoning drops exponentially:
$P(\text{correct}|C)\leq e^{-\delta(\text{TGC}(C)-\tau)}$ (12)
The proof is provided in Appendix A.2.
5.1 Transformation Generalization
Transformation generalization evaluates the ability of CoT reasoning to effectively transfer when models encounter novel transformations during testing, which is an especially prevalent scenario in real-world applications.
Figure 3: Performance of CoT reasoning on transformation generalization. Efficacy of CoT reasoning declines as the degree of distributional discrepancy increases.
Experimental Setup. To systematically evaluate the impact of transformations, we conduct experiments by varying transformations between training and testing sets while keeping other factors constant (e.g., elements, length, and format). Guided by the intuition formalized in Proposition 5.1, we define four incremental levels of distribution shift in transformations as shown in Figure 2: (i) In-Distribution (ID): The transformations in the test set are identical to those observed during training, e.g., $f_{1}\circ f_{1}\rightarrow f_{1}\circ f_{1}$. (ii) Composition (CMP): Test samples comprise novel compositions of previously encountered transformations, though each individual transformation remains familiar, e.g., ${f_{1}\circ f_{1},f_{1}\circ f_{2},f_{2}\circ f_{1}}\rightarrow f_{2}\circ f_{2}$. (iii) Partial Out-of-Distribution (POOD): Test data include compositions involving at least one novel transformation not seen during training, e.g., $f_{1}\circ f_{1}\rightarrow f_{1}\circ f_{2}$. (iv) Out-of-Distribution (OOD): The test set contains entirely novel transformation types that are unseen in training, e.g., $f_{1}\circ f_{1}\rightarrow f_{2}\circ f_{2}$.
Table 1: Full chain evaluation under different scenarios for transformation generalization.
Transformation (Train $\rightarrow$ Test) Scenario Exact Match Edit Distance BLEU Score
$f_{1}\circ f_{1}\rightarrow f_{1}\circ f_{1}$ ID 100.00% 0 1
$\{f_{2}\circ f_{2},f_{1}\circ f_{2},f_{2}\circ f_{1}\}\rightarrow f_{1}\circ f_{1}$ CMP 0.01% 0.1326 0.6867
$f_{1}\circ f_{2}\rightarrow f_{1}\circ f_{1}$ POOD 0.00% 0.1671 0.4538
$f_{2}\circ f_{2}\rightarrow f_{1}\circ f_{1}$ OOD 0.00% 0.2997 0.2947
Findings. Figure 3 illustrates the performance of the full chain under different distribution discrepancies computed by task generalize complexities (normalized between 0 and 1) in Definition 5.1. We can observe that, in general, the effectiveness of CoT reasoning decreases when distribution discrepancy increases. For the instance shown in Table 1, from in-distribution to composition, POOD, and OOD, the exact match decreases from 1 to 0.01, 0, and 0, and the edit distance increases from 0 to 0.13, 0.17 when tested on data with transformation $f_{1}\circ f_{1}$. Apart from ID, LLMs cannot produce a correct full chain in most cases, while they can produce correct CoT reasoning when exposed to some composition and POOD conditions by accident. As shown in Table 2, from $f_{1}\circ f_{2}$ to $f_{2}\circ f_{2}$, the LLMs can correctly answer 0.1% of questions. A close examination reveals that it is a coincidence, e.g., the query element is A, N, A, N, which happened to produce the same result for the two operations detailed in the Appendix D.1. When further analysis is performed by breaking the full chain into reasoning steps and answers, we observe strong consistency between the reasoning steps and answers. For example, under the composition generalization setting, the reasoning steps are entirely correct on test data distribution $f_{1}\circ f_{1}$ and $f_{2}\circ f_{2}$, but with wrong answers. Probe these insistent cases in Appendix D.1, we can find that when a novel transformation (say $f_{1}\circ f_{1}$) is present, LLMs try to generalize the reasoning paths based on the most similar ones (i.e., $f_{1}\circ f_{2}$) seen during training, which leads to correct reasoning paths, yet incorrect answer, which echo the example in the introduction. Similarly, generalization from $f_{1}\circ f_{2}$ to $f_{2}\circ f_{1}$ or vice versa allows LLMs to produce correct answers that are attributed to the commutative property between the two orthogonal transformations with unfaithful reasoning paths. Collectively, the above results indicate that the CoT reasoning fails to generalize to novel transformations, not even to novel composition transforms. Rather than demonstrating a true understanding of text, CoT reasoning under task transformations appears to reflect a replication of patterns learned during training.
Table 2: Evaluation on different components in CoT reasoning on transformation generalization. CoT reasoning shows inconsistency with the reasoning steps and answers.
Transformation (Train $\rightarrow$ Test) Exact Match Edit Distance
Reason Answer Full Chain Reason Answer Full Chain
$\{f_{1}\circ f_{1},f_{1}\circ f_{2},f_{2}\circ f_{1}\}\rightarrow f_{2}\circ f_{2}$ 100.00% 0.01% 0.01% 0.000 0.481 0.133
$\{f_{1}\circ f_{2},f_{2}\circ f_{1},f_{2}\circ f_{2}\}\rightarrow f_{1}\circ f_{1}$ 100.00% 0.01% 0.01% 0.000 0.481 0.133
$f_{1}\circ f_{2}\rightarrow f_{2}\circ f_{1}$ 0.00% 100.00% 0.00% 0.373 0.000 0.167
$f_{2}\circ f_{1}\rightarrow f_{1}\circ f_{2}$ 0.00% 100.00% 0.00% 0.373 0.000 0.167
Figure 4: Performance on unseen transformation using SFT in various levels of distribution shift. Introducing a small amount of unseen data helps CoT reasoning to generalize across different scenarios.
Experiment settings. To further probe when CoT reasoning can generalize to unseen transformations, we conduct supervised fine-tuning (SFT) on a small portion $\lambda$ of unseen data. In this way, we can decrease the distribution discrepancy between the training and test sets, which might help LLMs to generalize to test queries.
Findings. As shown in Figure 4, we can find that generally a very small portion ($\lambda=1.5e^{-4}$) of data can make the model quickly generalize to unseen transformations. The less discrepancy between the training and testing data, the quicker the model can generalize. This indicates that a similar pattern appears in the training data, helping LLMs to generalize to the test dataset.
5.2 Element Generalization
Element generalization is another critical factor to consider when LLMs try to generalize to new tasks.
Experiment settings. Similar to transformation generalization, we fix other factors and consider three progressive distribution shifts for elements: ID, CMP, and OOD, as shown in Figure 2. It is noted that in composition, we test if CoT reasoning can be generalized to novel combinations when seeing all the basic atoms in the elements, e.g., $(\texttt{A},\texttt{B},\texttt{C},\texttt{D})\rightarrow(\texttt{B},\texttt{C},\texttt{D},\texttt{A})$. Based on the atom order in combination (can be measured by edit distance $n$), the CMP can be further developed. While for OOD, atoms that constitute the elements are totally unseen during the training.
Figure 5: Element generalization results on various scenarios and relations.
Findings. Similar to transformation generalization, the performances degrade sharply when facing the distribution shift consistently across all transformations, as shown in Figure 5. From ID to CMP and OOD, the exact match decreases from 1.0 to 0 and 0, for all cases. Most strikingly, the BLEU score is 0 when transferred to $f_{1}$ and $f_{2}$ transformations. A failure case in Appendix D.1 shows that the models cannot respond to any words when novel elements are present. We further explore when CoT reasoning can generalize to novel elements by conducting SFT. The results are summarized in Figure 6. We evaluate the performance under three exact matches for the full chain under three scenarios, CMP based on the edit distance n. The result is similar to SFT on transformation. The performance increases rapidly when presented with similar (a small $n$) examples in the training data. Interestingly, the exact match rate for CoT reasoning aligns with the lower bound of performance when $n=3$, which might suggest the generalization of CoT reasoning on novel elements is very limited, even SFT on the downstream task. When we further analyze the exact match of reasoning, answer, and token during the training for $n=3$, as summarized in Figure 6b. We find that there is a mismatch of accuracy between the answer and the reasoning step during the training process, which somehow might provide an explanation regarding why CoT reasoning is inconsistent in some cases.
(a) Performance on unseen element via SFT in various CMP scenarios.
(b) Evaluation of CoT reasoning in SFT.
Figure 6: SFT performances for element generalization. SFT helps to generalize to novel elements.
6 Length Generalization
Length generalization examines how CoT reasoning degrades when models encounter test cases that differ in length from their training distribution. The difference in length could be introduced from the text space or the reasoning space of the problem. Therefore, we decompose length generalization into two complementary aspects: text length generalization and reasoning step generalization. Guided by instinct, we first propose to measure the length discrepancy.
Length Extrapolation Bound. We establish a power-law relationship for length extrapolation:
Proposition 6.1 (Length Extrapolation Gaussian Degradation).
For a model trained on chain-of-thought sequences of fixed length $L_{\text{train}}$, the generalization error at test length $L$ follows a Gaussian distribution:
$\mathcal{E}(L)=\mathcal{E}_{0}+\left(1-\mathcal{E}_{0}\right)\cdot\left(1-\exp\left(-\frac{\left(L-L_{\text{train }}\right)^{2}}{2\sigma^{2}}\right)\right)$ (13)
where $\mathcal{E}_{0}$ is the in-distribution error at $L=L_{\text{train}}$, $\sigma$ is the length generalization width parameter, and $L$ is the test sequence length
The proof is provided in Appendix A.3.
Table 3: Evaluation for text length generalization.
Length Exact Match (%) Edit Distance BLEU Score
Full Chain Reason Answer Full Chain Reason Answer Full Chain Reason Answer
2 0.00% 0.00% 0.00% 0.3772 0.4969 0.5000 0.4214 0.1186 0.0000
3 0.00% 0.00% 0.00% 0.2221 0.3203 0.2540 0.5471 0.1519 0.0000
4 100.00% 100.00% 100.00% 0.0000 0.0000 0.0000 1.0000 1.0000 1.0000
5 0.00% 0.00% 0.00% 0.1818 0.2667 0.2000 0.6220 0.1958 0.2688
6 0.00% 0.00% 0.00% 0.3294 0.4816 0.3337 0.4763 0.1174 0.2077
6.1 Text Length Generalization
Text length generalization evaluates how CoT performance varies when the input text length (i.e., the element length $l$) differs from training examples. Considering the way LLMs process long text, this aspect is crucial because real-world problems often involve varying degrees of complexity that manifest as differences in problem statement length, context size, or information density.
Figure 7: Performance of text length generalization across various padding strategies. Group strategies contribute to length generalization.
Experiment settings. We pre-train LLMs on the dataset with text length merely on $l=4$ while fixing other factors and evaluate the performance on a variety of lengths. We consider three different padding strategies during the pre-training: (i) None: LLMs do not use any padding. (ii) Padding: We pad LLM to the max length of the context window. (iii) Group: We group the text and truncate it into segments with a maximum length.
Findings. As illustrated in the Table 3, the CoT reasoning failed to directly generate two test cases even though those lengths present a mild distribution shift. Further, the performance declines as the length discrepancy increases shown in Figure 7. For instance, from data with $l=4$ to those with $l=3$ or $l=5$, the BLEU score decreases from 1 to 0.55 and 0.62. Examples in Appendix D.1 indicate that LLMs attempt to produce CoT reasoning with the same length as the training data by adding or removing tokens in the reasoning chains. The efficacy of CoT reasoning length generalization deteriorates as the discrepancy increases. Moreover, we consider using a different padding strategy to decrease the divergence between the training data and test cases. We found that padding to the max length doesn’t contribute to length generalization. However, the performance increases when we replace the padding with text by using the group strategy, which indicates its effectiveness.
6.2 Reasoning Step Generalization
The reasoning step generalization investigates whether models can extrapolate to reasoning chains requiring different steps $k$ from those observed during training. which is a popular setting in multi-step reasoning tasks.
Experiment settings. Similar to text length generalization, we first pre-train the LLM with reasoning step $k=2$, and evaluate on data with reasoning step $k=1$ or $k=3$.
(a) Reasoning step. From k=2 to k=1
(b) Reasoning step. From k=2 to k=3
Figure 8: Test performance for reasoning-step generalization across varying training data compositions. Performance varies systematically with changes in the distribution of training data components.
Findings. As showcased in Figure 8, CoT reasoning cannot generalize across data requiring different reasoning steps, indicating the failure of generalization. Then, we try to decrease the distribution discrepancy introduced by gradually increasing the ratio of unseen data while keeping the dataset size the same when pre-training the model. And then, we evaluate the performance on two datasets. As we can observe, the performance on the target dataset increases along with the ratio. At the same time, the LLMs can not generalize to the original training dataset because of the small amount of training data. The trend is similar when testing different-step generalization, which follows the intuition and validates our hypothesis directly.
7 Format Generalization
Format generalization assesses the robustness of CoT reasoning to surface-level variations in test queries. This dimension is especially crucial for determining whether models have internalized flexible, transferable reasoning strategies or remain reliant on the specific templates and phrasings encountered during training.
Format Alignment Score. We introduce a metric for measuring prompt similarity:
Definition 7.1 (Format Alignment Score).
For training prompt distribution $P_{train}$ and test prompt $p_{test}$:
$\text{PAS}(p_{test})=\max_{p\in P_{train}}\cos(\phi(p),\phi(p_{test}))$ (14)
where $\phi$ is a prompt embedding function.
(a) Format generalization. Performance under various perturbation methods.
(b) Format generalization. Performance vs. various applied perturbation areas.
Figure 9: Performance of format generalization. Testing performance varies with different noise levels and areas where the noise is applied.
Experiment settings. To systematically probe this, we introduce four distinct perturbation modes to simulate scenario in real-world: (i) insertion, where a noise token is inserted before each original token; (ii) deletion: it deletes the original token; (iii) modification: it replaces the original token with a noise token; and (iv) hybrid mode: it combines multiple perturbations. Each mode is applied for tokens with probabilities $p$, enabling us to quantify the model’s resilience to increasing degrees of prompt distribution shift.
Findings. As shown in Figure 9a, we found that generally CoT reasoning can be easily affected by the format changes. No matter insertion, deletion, modifications, or hybrid mode, it creates a format discrepancy that affects the correctness. Among them, the deletion slightly affects the performance. While the insertions are relatively highly influential on the results. We further divide the query into several sections: elements, transformations, and prompt tokens. As shown in Figure 9b, we found that the elements and transformation play an important role in the format, whereas the changes to other tokens rarely affect the results.
8 Temperature and Model Size
Temperature and model size generalization explores how variations in sampling temperature and model capacity can influence the stability and robustness of CoT reasoning. For the sake of rigorous evaluation, we further investigate whether different choices of temperatures and model sizes may significantly affect our results.
Experiment settings. We explore the impact of different temperatures on the validity of the presented results. We adopt the same setting in the transformation generalization.
Findings. As illustrated in Figure 10a, LLMs tend to generate consistent and reliable CoT reasoning across a broad range of temperature settings (e.g., from 1e-5 up to 1), provided the values remain within a suitable range. This stability is maintained even when the models are evaluated under a variety of distribution shifts.
(a) Influences of various temperatures.
(b) Influences of various sizes.
Figure 10: Temperature and model size. The findings hold under different temperatures and model sizes.
Experiment settings. We further examine the influence of model size by employing the same experimental configuration as used in the novel relation SFT study. In particular, we first pretrain models of different sizes using the transformation $f_{1}\circ f_{1}$, and subsequently perform SFT on $f_{2}\circ f_{2}$ while varying the SFT ratios.
Finding. Fig. 10b shows the accuracy of models with different sizes using different SFT ratios, which closely matches the result of our default model size across all evaluated settings and configurations.
9 Discussion and Implication
Our investigation, conducted through the controlled environment of DataAlchemy, reveals that the apparent reasoning prowess of Chain-of-Thought (CoT) is largely a brittle mirage. The findings across task, length, and format generalization experiments converge on a conclusion: CoT is not a mechanism for genuine logical inference but rather a sophisticated form of structured pattern matching, fundamentally bounded by the data distribution seen during training. When pushed even slightly beyond this distribution, its performance degrades significantly, exposing the superficial nature of the “reasoning” it produces.
While our experiments utilized models trained from scratch in a controlled environment, the principles uncovered are extensible to large-scale pre-trained models. We summarize the implications for practitioners as follows.
Guard Against Over-reliance and False Confidence. CoT should not be treated as a “plug-and-play” module for robust reasoning, especially in high-stakes domains like medicine, finance, or legal analysis. The ability of LLMs to produce “fluent nonsense”—plausible but logically flawed reasoning chains—can be more deceptive and damaging than an outright incorrect answer, as it projects a false aura of dependability. Sufficient auditing from domain experts is indispensable.
Prioritize Out-of-Distribution (OOD) Testing. Standard validation practices, where the test set closely mirrors the training set, are insufficient to gauge the true robustness of a CoT-enabled system. Practitioners must implement rigorous adversarial and OOD testing that systematically probes for vulnerabilities across task, length, and format variations.
Recognize Fine-Tuning as a Patch, Not a Panacea. Our results show that Supervised Fine-Tuning (SFT) can quickly “patch” a model’s performance on a new, specific data distribution. However, this should not be mistaken for achieving true generalization. It simply expands the model’s “in-distribution” bubble slightly. Relying on SFT to fix every OOD failure is an unsustainable and reactive strategy that fails to address the core issue: the model’s lack of abstract reasoning capability.
10 Conclusion
In this paper, we critically examine the COT reasoning of LLMs through the lens of data distribution, revealing that the perceived structured reasoning capability largely arises from inductive biases shaped by in-distribution training data. We propose a controlled environment, DataAlchemy, allowing systematic probing of CoT reasoning along three crucial dimensions: task structure, reasoning length, and query format. Empirical findings consistently demonstrate that CoT reasoning effectively reproduces reasoning patterns closely aligned with training distributions but suffers significant degradation when faced with distributional deviations. Such observations reveal the inherent brittleness and superficiality of current CoT reasoning capabilities. We provide insights that emphasize real-world implications for both practitioners and researchers.
[…]
Appendix A Proof of Theorems
A.1 Proof of CoT Generalization Bound
Proof.
Let $f_{\theta}$ be a model trained on samples from the distribution $\mathcal{D}_{\text{train}}$ using a loss function $\ell(f_{\theta}(x),y)$ that is $\Lambda$-Lipschitz and bounded. The expected test risk is given by
$R_{\text{test}}(f_{\theta})=\mathbb{E}_{(x,y)\sim\mathcal{D}_{\text{test}}}\left[\ell(f_{\theta}(x),y)\right].$ (15)
We can decompose the test risk as
$R_{\text{test}}(f_{\theta})=R_{\text{train}}(f_{\theta})+\left(R_{\text{test}}(f_{\theta})-R_{\text{train}}(f_{\theta})\right).$ (16)
To bound the discrepancy between $R_{\text{test}}$ and $R_{\text{train}}$, we invoke a standard result from statistical learning theory. Given that $\ell$ is $\Lambda$-Lipschitz and the discrepancy measure $\Delta(\mathcal{D}_{\text{train}},\mathcal{D}_{\text{test}})$ is an integral probability metric (e.g., Wasserstein-1 distance), we have
$\left|R_{\text{test}}(f_{\theta})-R_{\text{train}}(f_{\theta})\right|\leq\Lambda\cdot\Delta(\mathcal{D}_{\text{train}},\mathcal{D}_{\text{test}}).$ (17)
Therefore, the test risk satisfies
$R_{\text{test}}(f_{\theta})\leq R_{\text{train}}(f_{\theta})+\Lambda\cdot\Delta(\mathcal{D}_{\text{train}},\mathcal{D}_{\text{test}}).$ (18)
We next account for the generalization gap between the empirical training risk $\hat{R}_{\text{train}}(f_{\theta})$ and the expected training risk $R_{\text{train}}(f_{\theta})$. By applying a concentration inequality (e.g., Hoeffding’s inequality), with probability at least $1-\delta$, we have
$R_{\text{train}}(f_{\theta})\leq\hat{R}_{\text{train}}(f_{\theta})+\mathcal{O}\left(\sqrt{\frac{\log(1/\delta)}{n}}\right),$ (19)
where $n$ is the number of training samples.
Combining the above, we obtain that with high probability,
$R_{\text{test}}(f_{\theta})\leq\hat{R}_{\text{train}}(f_{\theta})+\Lambda\cdot\Delta(\mathcal{D}_{\text{train}},\mathcal{D}_{\text{test}})+\mathcal{O}\left(\sqrt{\frac{\log(1/\delta)}{n}}\right).$ (20)
This concludes the proof. ∎
A.2 Proof of Task Generalization Failure Threshold
We establish the exponential decay bound through a probabilistic analysis of reasoning failure modes in the presence of task generalization complexity.
Let $\Omega$ denote the sample space of all possible reasoning configurations, and let $C\in\Omega$ represent a specific configuration. We define the following events: $A_{i}$ as the event that element $a_{i}$ is novel, i.e., $a_{i}\notin\mathcal{E}^{i}_{\text{train}}$; $F_{j}$ as the event that transformation $f_{j}$ is novel, i.e., $f_{j}\notin\mathcal{F}_{\text{train}}$; and $\mathcal{Q}$ as the event that the transformation sequence $(f_{1},f_{2},\ldots,f_{k})$ is novel, i.e., $(f_{1},f_{2},\ldots,f_{k})\notin\mathcal{P}_{\text{train}}$.
Here we make the assumption that the reasoning failures induced by novel arguments, functions, and patterns contribute independently to the overall failure probability and hence we model the success probability as a product of component-wise success rates:
$P(\text{correct}|C)=P_{0}\prod_{i=1}^{m}\rho_{a}^{\mathbb{I}[A_{i}]}\prod_{j=1}^{n}\rho_{f}^{\mathbb{I}[F_{j}]}\rho_{p}^{\mathbb{I}[\mathcal{Q}]}\rho_{c}^{C_{T}}$
where $P_{0}\in(0,1]$ represents the baseline success probability when all components are within the training distribution, and $\rho_{a},\rho_{f},\rho_{p},\rho_{c}\in(0,1)$ are the degradation factors associated with novel arguments, functions, patterns, and task-specific complexity, respectively.
$\ln P(\text{correct}\mid C)\;=$ $\ln P_{0}+\sum_{i=1}^{m}\mathbb{I}[A_{i}]\,\ln\rho_{a}+\sum_{j=1}^{n}\mathbb{I}[F_{j}]\,\ln\rho_{f}+\mathbb{I}[\mathcal{Q}]\,\ln\rho_{p}+C_{T}\,\ln\rho_{c}$ (21)
For notational convenience, we define the positive constants:
$\xi_{a}:=-\ln\rho_{a}>0,\xi_{f}:=-\ln\rho_{f}>0,\xi_{p}:=-\ln\rho_{p}>0,\xi_{c}:=-\ln\rho_{c}>0$
hence we have:
$\ln P(\text{correct}|C)=\ln P_{0}-\xi_{a}\sum_{i=1}^{m}\mathbb{I}[A_{i}]-\xi_{f}\sum_{j=1}^{n}\mathbb{I}[F_{j}]-\xi_{p}\mathbb{I}[\mathcal{Q}]-\xi_{c}C_{T}$ (22)
Lemma: Relationship to TGC. The expression in equation above can be bounded in terms of $\text{TGC}(C)$ as follows:
$\ln P(\text{correct}|C)\leq\ln P_{0}-\delta\cdot\text{TGC}(C)$ (23)
where $\delta=\min(\frac{\xi_{a}}{\alpha},\frac{\xi_{f}}{\beta},\frac{\xi_{p}}{\gamma},\xi_{c})>0$.
Proof of Lemma: From the definition of $\text{TGC}(C)$ in Eq. (11), we have:
$\text{TGC}(C)=\alpha\sum_{i=1}^{m}\mathbb{I}[A_{i}]+\beta\sum_{j=1}^{n}\mathbb{I}[F_{j}]+\gamma\mathbb{I}[\mathcal{Q}]+C_{T}$ (24)
By the definition of $\delta$, each term in Eq. (22) satisfies:
$\xi_{a}\sum_{i=1}^{m}\mathbb{I}[A_{i}]\geq\delta\alpha\sum_{i=1}^{m}\mathbb{I}[A_{i}]$ (25)
$\xi_{f}\sum_{j=1}^{n}\mathbb{I}[F_{j}]\geq\delta\beta\sum_{j=1}^{n}\mathbb{I}[F_{j}]$ (26)
$\xi_{p}\mathbb{I}[\mathcal{Q}]\geq\delta\gamma\mathbb{I}[\mathcal{Q}]$ (27)
$\xi_{c}C_{T}\geq\delta C_{T}$ (28)
Summing these inequalities establishes Eq. (23).
We now define the threshold $\tau:=\frac{\ln P_{0}}{\delta}$. From Eq. (23), when $\text{TGC}(C)>\tau$, we have:
$\ln P(\operatorname{correct}\mid C)$ $\leq\ln P_{0}-\delta\cdot\operatorname{TGC}(C)$ (29)
$=\delta(\tau-\operatorname{TGC}(C))$ (30)
$=-\delta(\operatorname{TGC}(C)-\tau)$ (31)
Exponentiating both sides yields the desired bound: $P(\operatorname{correct}\mid C)\leq e^{-\delta(\mathrm{TGC}(C)-\tau)}$
A.3 Proof of Length Extrapolation Bound
Proof.
Consider a transformer model $f_{\theta}$ processing sequences of length $L$. The model implicitly learns position-dependent representations through positional encodings $\text{PE}(i)\in\mathbb{R}^{d}$ for position $i\in\{1,\ldots,L\}$ and attention patterns $A_{ij}=\text{softmax}\left(\frac{Q_{i}K_{j}^{T}}{\sqrt{d}}\right)$.
During training on fixed length $L_{\text{train}}$, the model learns a specific distribution:
$p_{\text{train}}(\mathbf{h})=p(\mathbf{h}\mid L=L_{\text{train}})$ (32)
where $\mathbf{h}=\{h_{1},\ldots,h_{L}\}$ represents hidden states.
For sequences of length $L\neq L_{\text{train}}$, we encounter distribution shift in two forms: (1) positional encoding mismatch, where the model has never seen positions $i>L_{\text{train}}$ if $L>L_{\text{train}}$, and (2) attention pattern disruption, where the learned attention patterns are calibrated for length $L_{\text{train}}$.
The KL divergence between training and test distributions can be bounded:
$D_{KL}(p_{\text{test}}\|p_{\text{train}})\propto|L-L_{\text{train}}|^{2}$ (33)
This quadratic relationship arises from linear accumulation of positional encoding errors and quadratic growth in attention pattern misalignment due to pairwise interactions.
Let $\mathcal{E}(L)$ be the prediction error at length $L$. We decompose it as:
$\mathcal{E}(L)=\mathcal{E}_{\text{inherent}}(L)+\mathcal{E}_{\text{shift}}(L)$ (34)
where $\mathcal{E}_{\text{inherent}}(L)=\mathcal{E}_{0}$ is the inherent model error (constant) and $\mathcal{E}_{\text{shift}}(L)$ is the error due to distribution shift.
The distribution shift error follows from the Central Limit Theorem. As the error accumulates over sequence positions, the total shift error converges to:
$\mathcal{E}_{\text{shift}}(L)=(1-\mathcal{E}_{0})\cdot\left(1-\exp\left(-\frac{(L-L_{\text{train}})^{2}}{2\sigma^{2}}\right)\right)$ (35)
This form ensures that $\mathcal{E}_{\text{shift}}(L_{\text{train}})=0$ (no shift at training length) and $\lim_{|L-L_{\text{train}}|\to\infty}\mathcal{E}_{\text{shift}}(L)=1-\mathcal{E}_{0}$ (maximum error bounded by 1).
The width parameter $\sigma$ depends on:
$\sigma=\sigma_{0}\cdot\sqrt{\frac{d}{L_{\text{train}}}}$ (36)
where $\sigma_{0}$ is a model-specific constant, $d$ is the model dimension, and the $\sqrt{d/L_{\text{train}}}$ factor captures the concentration of measure in high dimensions.
Therefore, the total error follows:
$\mathcal{E}(L)=\mathcal{E}_{0}+(1-\mathcal{E}_{0})\cdot\left(1-\exp\left(-\frac{(L-L_{\text{train}})^{2}}{2\sigma^{2}}\right)\right)$ (37)
This Gaussian form naturally emerges from the accumulation of position-dependent errors and matches the experimental observation of near-zero error at $L=L_{\text{train}}$ with symmetric increase in both directions. ∎
Appendix B Experiment Details
The model is trained using the AdamW optimiser in mixed precision (FP16). The default learning rate is $3\times 10^{-3}$, and the schedule follows a cosine decay with a 10% warm-up ratio. Training is conducted for 10 epochs, using a batch size of 1024. A weight decay of 0.01 is applied, and gradient norms are clipped at 1.0.
Appendix C Illustration of Datasets
Below are the examples of transformation $f_{1}$ and $f_{2}$:
Transformation[F1]: A A F Q [F1] <answer> N N S D
Transformation[F2]: A A L P [F2] <answer> A L P A
aside from single transformation, we can composite transformations arbitrarily:
Transformation[F1F2]: A C I A [F1] [F2] <think>
N P V N [F2] <answer>
P V N N
Transformation[$F2$]: N O V S [F2] [F2] <think>
O V S N [F2] <answer>
V S N O
we use exact match, edit distance and BELU score to measure the discrepancy between generated tokens and the labels. For more than one transformation examples, we can further measure the discrepancy for reasoning and answering seperately.
Appendix D Additional Experimental Results
D.1 Additional Qualitative Analysis
D.1.1 Orthognal Transformation Caused Coincidence
The following case shows that even if the transformation is different, the model that trained on transformation $f_{2}\circ f_{1}$ can still provide correct answer through incorrect reasoning:
Prompt: A A A B [F1] [F2] <answer>’
Generated:’B A A A [F1] <answer> O N N N
Expected: O N N N
D.1.2 Correct reasoning but failed in final answer
The following case shows that the model pretrained on the union of three transformation $f_{1}\circ f_{2},f_{2}\circ f_{1},f_{2}\circ f_{2}$ and test on $f_{1}\circ f_{1}$
Prompt: A A A D [R1] [R1] <think>’
Generated: N N N Q [R1] <answer> N N Q N
Expected: N N N Q [R1] <answer> A A A D
D.1.3 Failure to generalize to novel element
The following case shows that the model trained on element set $a_{i}\in[A,M]$ can not generalize to unseen elements such as N or O
Prompt: N N N O [F1] [F1] <think>’
Generated: R V Q S [F1] <answer> E I D F
Expected: A A A B [F1] <answer> N N N O
D.1.4 LLM reproduces CoT reasoning at seen lengths
The following case shows that model trained under $f_{1}\circ f_{1}$ tried to reproduce the length in training data by adding tokens in the reason chain even prompted with seen transformation $f_{1}$
Prompt: A A B D [f1] <answer>’
Generated: N O A Z N N O Q [f1]
<answer> A A B D
Expected: N N O Q

### v4

@@
 Abstract
-Chain-of-Thought (CoT) prompting has been shown to improve Large Language Model (LLM) performance on various tasks.
-With this approach, LLMs appear to produce human-like reasoning steps before providing answers (a.k.a., CoT reasoning), which often leads to the perception that they engage in deliberate inferential processes.
-However, some initial findings suggest that CoT reasoning may be more superficial than it appears, motivating us to explore further.
-In this paper, we study CoT reasoning via a data distribution lens and investigate if CoT reasoning reflects a structured inductive bias learned from in-distribution data, allowing the model to conditionally generate reasoning paths that approximate those seen during training.
-Thus, its effectiveness is fundamentally bounded by the degree of distribution discrepancy between the training data and the test queries.
-With this lens, we dissect CoT reasoning via three dimensions: task, length, and format.
-To investigate each dimension, we design DataAlchemy, an isolated and controlled environment to train LLMs from scratch and systematically probe them under various distribution conditions.
-Our results reveal that CoT reasoning is a brittle mirage that vanishes when it is pushed beyond training distributions.
-This work offers a deeper understanding of why and when CoT reasoning fails, emphasizing the ongoing challenge of achieving genuine and generalizable reasoning.
+Chain-of-Thought (CoT) prompting has been shown to be effective in eliciting structured reasoning (i.e., CoT reasoning) from large language models (LLMs).
+Regardless of its popularity, recent studies expose its failures in some reasoning tasks, raising fundamental questions about the nature of CoT reasoning.
+In this work, we propose a data distribution lens to understand when and why CoT reasoning succeeds or fails.
+We hypothesize that CoT reasoning reflects a structured inductive bias learned from in-distribution data, enabling models to conditionally generate reasoning trajectories that approximate those observed during training.
+As such, the effectiveness of CoT reasoning is fundamentally governed by the nature and degree of distribution discrepancy between training data and test queries.
+Guided by this lens, we dissect CoT reasoning via three dimensions: task, length, and format.
+To test the hypothesis, we introduce DataAlchemy, an abstract and fully controllable environment that trains LLMs from scratch and systematically probes them under various distribution conditions.
+Through rigorous controlled experiments, we reveal that CoT reasoning is a brittle mirage when it is pushed beyond training distributions, emphasizing the ongoing challenge of achieving genuine and generalizable reasoning.
 Our code is available at GitHub: https://github.com/ChengshuaiZhao0/DataAlchemy.
@@
 Figure 1: The data perspective lens.
-CoT reasoning’s effectiveness is fundamentally bounded by the degree of distribution discrepancy between the training data and the test queries.
-Guided by this lens, we dissect CoT reasoning via three dimensions: task, length, and format.
-Recent years have witnessed Large Language Models’ (LLMs) dominant role in various domains (Zhao et al., 2023; Li et al., 2025b; Zhao et al., 2025; Ting et al., 2025) through versatile prompting techniques (Wei et al., 2022; Yao et al., 2023; Kojima et al., 2022).
-Among these, Chain-of-Thought (CoT) prompting (Wei et al., 2022) has emerged as a prominent method for eliciting structured reasoning from LLMs (a.k.a., CoT reasoning).
-By appending a simple cue such as “Let’s think step by step,” LLMs decompose complex problems into intermediate steps, producing outputs that resemble human-like reasoning.
-It has been shown to be effective in tasks requiring logical inference(Xu et al., 2024), mathematical problem solving (Imani et al., 2023), and commonsense reasoning (Wei et al., 2022).
-The empirical successes of CoT reasoning lead to the perception that LLMs engage in deliberate inferential processes (Yu et al., 2023; Zhang et al., 2024a; Ling et al., 2023; Zhang et al., 2024c).
-However, a closer examination reveals inconsistencies that challenge this optimistic view.
-Consider this straightforward question: “The day the US was established is in a leap year or a normal year?” When prompted with the CoT prefix, the modern LLM Gemini responded: “The United States was established in 1776.
-1776 is divisible by 4, but it’s not a century year, so it’s a leap year.
-Therefore, the day the US was established was in a normal year.” This response exemplifies a concerning pattern: the model correctly recites the leap year rule and articulates intermediate reasoning steps, yet produces a logically inconsistent conclusion (i.e., asserting 1776 is both a leap year and a normal year).
-Such inconsistencies suggest that there is a distinction between human-like inference and CoT reasoning.
-An expanding body of analyses reveals that LLMs tend to rely on surface-level semantics and clues rather than logical procedures (Stechly et al., 2024; Kambhampati, 2024; Chen et al., 2025b; Lanham et al., 2023).
-LLMs construct superficial chains of logic based on learned token associations, often failing on tasks that deviate from commonsense heuristics or familiar templates (Tang et al., 2023).
-In the reasoning process, performance degrades sharply when irrelevant clauses are introduced, which indicates that models cannot grasp the underlying logic (Mirzadeh et al., 2024).
-This fragility becomes even more apparent when models are tested on more complex tasks, where they frequently produce incoherent solutions and fail to follow consistent reasoning paths (Shojaee et al., 2025).
-Collectively, these pioneering works deepen the skepticism surrounding the true nature of CoT reasoning.
-In light of this line of research, we question the CoT reasoning by proposing an alternative lens through data distribution and further investigating why and when it fails.
-We hypothesize that CoT reasoning reflects a structured inductive bias learned from in-distribution data, allowing the model to conditionally generate reasoning paths that approximate those seen during training.
-As such, its effectiveness is inherently limited by the nature and extent of the distribution discrepancy between training data and the test queries.
-Guided by this data distribution lens, we dissect CoT reasoning via three dimensions: (i) task—To what extent CoT reasoning can handle tasks that involve transformations or previously unseen task structures.
-(2) length—how CoT reasoning generalizes to chains with length different from that of training data; and (3) format—how sensitive CoT reasoning is to surface-level query form variations.
-To evaluate each aspect, we introduce DataAlchemy, a controlled and isolated experiment that allows us to train LLMs from scratch and systematically probe them under various distribution shifts.
-Our findings reveal that CoT reasoning works effectively when applied to in-distribution or near in-distribution data but becomes fragile and prone to failure even under moderate distribution shifts.
+CoT reasoning’s effectiveness is fundamentally bounded by the degree of distribution discrepancy introduced by task, length, and format between the training data and the test queries.
+Chain-of-Thought (CoT) prompting wei2022chain has emerged as a prominent method for eliciting structured reasoning from LLMs (a.k.a., CoT reasoning).
+By appending a simple cue such as “Let’s think step by step”, LLMs decompose complex problems into intermediate steps, producing outputs that resemble human-like reasoning.
+It has been shown to be effective in tasks requiring logical inference xu2024faithful, mathematical problem solving imani2023mathprompter, and commonsense reasoning wei2022chain.
+The empirical success led to CoT reasoning being seen as a promising direction towards artificial general intelligence.
+However, some pioneering studies have revealed failures that challenge this optimistic view mirzadehgsm.
+stechly2024chain demonstrate that LLMs fail to generalize in planning tasks, revealing a deficiency in true algorithmic reasoning.
+shojaee2025illusion find that reasoning models experience an accuracy collapse in puzzle-solving once task complexity exceeds a critical threshold.
+sun2025omega demonstrate that LLMs struggle to solve complex mathematical reasoning problem, failing to integrate or adapt learned skills to novel or creative tasks.
+Considering the above opposing opinions, there clearly lacks an indisputable lens to understand why and when CoT reasoning succeeds or fails.
+Current evaluation approaches have intrinsic limitations that prevent them from answering the Why and When questions: (i) Narrowly defined settings.
+Existing frameworks focus on specific tasks and evaluate them using specific LLMs, thereby overlooking common structural patterns and characteristics.
+(ii) Data entanglement.
+Most evaluations are conducted in real-world scenarios, where the complexity precludes fully controlled experiments to isolate fine-grained factors.
+(iii) Data leakage.
+Pre-trained LLMs suffer from data leakage and benchmark containment problems, undermining the effectiveness and validity of evaluations.
+In this work, we study CoT reasoning by introducing a data distribution lens.
+Specifically, we hypothesize that CoT reasoning reflects a structured inductive bias learned from in-distribution data, enabling models to conditionally generate reasoning trajectories that approximate those observed during training.
+As such, the effectiveness of CoT reasoning is fundamentally governed by the nature and degree of distribution discrepancy between training data and test queries.
+Guided by this lens, we revisit existing NLP tasks and identify three primary axes along which distribution shifts may occur: task (i.e., unseen task structures), length (i.e., different text lengths and reasoning lengths), and format (i.e., query format variant).
+To tackle the issue of evaluations and validate our hypothesis, we further introduce DataAlchemy, an abstract, controllable, and clean environment.
+DataAlchemy provides an abstract representation system that distills various real-world NLP tasks into key components: atoms (i.e., token space), elements (i.e., text space), and transformations (i.e., operation space).
+By varying these components, we curate data that exhibits various distribution discrepancies, naturally achieving full and fine-grained control over the entire evaluation pipeline.
+Later, we train models from scratch to avoid data leakage and employ controlled experiments to rigorously test our hypotheses.
+Our findings reveal that CoT reasoning works effectively when applied to (near) in-distribution data, but becomes fragile and prone to failure even under moderate distribution shifts.
 In some cases, LLMs generate fluent yet logically inconsistent reasoning steps.
 The results suggest that what appears to be structured reasoning can be a mirage, emerging from memorized or interpolated patterns in the training data rather than logical inference.
-These insights carry important implications for both practitioners and researchers.
-For practitioners, our results highlight the risk of relying on CoT as a plug-and-play solution for reasoning tasks and caution against equating CoT-style output with human thinking.
-For researchers, the results underscore the ongoing challenge of achieving reasoning that is both faithful and generalizable, motivating the need to develop models that can move beyond surface-level pattern recognition to exhibit deeper inferential competence.
-Our contributions are summarized as follows:
+Our contributions can be summarized as follows:
   • $\star$
     Novel perspective.
-We propose a data distribution lens for CoT reasoning, illuminating that its effectiveness stems from structured inductive biases learned from in-distribution training data.
-This framework provides a principled lens for understanding why and when CoT reasoning succeeds or fails.
+We propose a data distribution lens for CoT reasoning, revealing that its effectiveness arises from structured inductive biases learned from in-distribution data.
+This lens offers a principled foundation for understanding why and when CoT reasoning succeeds or fails.
   • $\star$
-    Controlled environment.
-We introduce DataAlchemy, an isolated experimental framework that enables training LLMs from scratch and systematically probing CoT reasoning.
-This controlled setting allows us to isolate and analyze the effects of distribution shifts on CoT reasoning without interference from complex patterns learned during large-scale pre-training.
+    Controllable environment.
+We develop an abstract, fully controllable, and clean environment—DataAlchemy that abstracts NLP tasks, enabling systematic analysis of CoT reasoning under distribution discrepancies.
+DataAlchemy can serve as a research platform for probing the intrinsic behavior of LLMs, and facilitating the discovery of scientific principles.
   • $\star$
-    Empirical validation.
-We conduct systematic empirical validation across three critical dimensions—task, length, and format.
-Our experiments demonstrate that CoT reasoning exhibits sharp performance degradation under distribution shifts, revealing that seemingly coherent reasoning masks shallow pattern replication.
+    Rigorous investigation.
+Guided by data distribution lens, we dissect the CoT reasoning via three dimensions: task, length, and format.
+Later, we curate data that reflects fine-grained factors in each dimension and conduct controlled experiments to isolate and examine each factor.
   • $\star$
-    Real-world implication.
-This work reframes the understanding of contemporary LLMs’ reasoning capabilities and emphasizes the risk of over-reliance on COT reasoning as a universal problem-solving paradigm.
-It underscores the necessity for proper evaluation methods and the development of LLMs that possess authentic and generalizable reasoning capabilities.
+    General validity.
+We train and fine-tune hundreds of LLMs with varying sizes (from 62K to 14B), architectures (e.g., GPT, LLaMA, and Qwen), and temperatures (from 1e-5 to 10).
+The results consistently show that the effectiveness of CoT reasoning varies with the degree of distribution discrepancy, substantiating the generality of the proposed data distribution lens.
 2 Related Work
 2.1 LLM Prompting and CoT
-Chain-of-Thought (CoT) prompting revolutionized how we elicit reasoning from Large Language Models by decomposing complex problems into intermediate steps (Wei et al., 2022).
-By augmenting few-shot exemplars with reasoning chains, CoT showed substantial performance gains on various tasks (Xu et al., 2024; Imani et al., 2023; Wei et al., 2022).
-Building on this, several variants emerged.
-Zero-shot CoT triggers reasoning without exemplars using instructional prompts (Kojima et al., 2022), and self-consistency enhances performance via majority voting over sampled chains (Wang et al., 2023).
-To reduce manual effort, Auto-CoT generates CoT exemplars using the models themselves (Zhang et al., 2023).
-Beyond linear chains, Tree-of-Thought (ToT) frames CoT as a tree search over partial reasoning paths (Yao et al., 2023), enabling lookahead and backtracking.
-SymbCoT combines symbolic reasoning with CoT by converting problems into formal representations (Xu et al., 2024).
-Recent work increasingly integrates CoT into the LLM inference process, generating long-form CoTs (Jaech et al., 2024; Team, 2024; Guo et al., 2025; Team et al., 2025).
-This enables flexible strategies like mistake correction, step decomposition, reflection, and alternative reasoning paths (Yeo et al., 2025; Chen et al., 2025a).
-The success of prompting techniques and long-form CoTs has led many to view them as evidence of emergent, human-like reasoning in LLMs.
-In this work, we challenge that viewpoint by adopting a data-centric perspective and demonstrating that CoT behavior arises largely from pattern matching over training distributions.
+Chain-of-Thought (CoT) prompting improves large language model performance by eliciting intermediate reasoning steps for complex problems (wei2022chain).
+Extensions include zero-shot CoT (kojima2022large), self-consistency via sampling and voting (wang2023selfconsistency), and Auto-CoT, which automatically generates reasoning exemplars (zhang2023automatic).
+Beyond linear reasoning, Tree-of-Thought enables search over multiple reasoning paths (yao2023tree), while SymbCoT integrates symbolic representations into CoT (xu2024faithful).
+More recent work embeds long-form CoT directly into inference, enabling reflection, error correction, and alternative reasoning strategies (jaech2024openai; team2024qwq; guo2025deepseek; team2025kimi; yeo2025demystifying; chen2025towards).
+In this work, we investigate whether CoT reflects genuine reasoning or merely pattern interpolation.
 2.2 Discussion on Illusion of LLM Reasoning
-While Chain-of-Thought prompting has led to impressive gains on complex reasoning tasks, a growing body of work has started questioning the nature of these gains (Stechly et al., 2024, 2025; Kambhampati et al., 2025).
-One major line of research highlights the fragility of CoT reasoning.
-Minor and semantically irrelevant perturbations such as distractor phrases or altered symbolic forms can cause significant performance drops in state-of-the-art models (Mirzadeh et al., 2024; Tang et al., 2023).
-Models often incorporate such irrelevant details into their reasoning, revealing a lack of sensitivity to salient information.
-Other studies show that models prioritize the surface form of reasoning over logical soundness; in some cases, longer but flawed reasoning paths yield better final answers than shorter, correct ones (Bentham et al., 2024).
-Similarly, performance does not scale with problem complexity as expected—models may overthink easy problems and give up on harder ones (Shojaee et al., 2025).
-Another critical concern is the faithfulness of the reasoning process.
-Intervention-based studies reveal that final answers often remain unchanged even when intermediate steps are falsified or omitted (Lanham et al., 2023), a phenomenon dubbed the illusion of transparency (Bentham et al., 2024; Chen et al., 2025b).
-Together, these findings suggest that LLMs are not principled reasoners but rather sophisticated simulators of reasoning-like text.
-However, a systematic understanding of why and when CoT reasoning fails is still a mystery.
+Recent work questions the robustness and faithfulness of these gains (stechly2024chain).
+A prominent line of research shows that CoT reasoning is highly fragile: semantically irrelevant perturbations, such as distractor phrases or altered symbolic representations, can substantially degrade performance (mirzadehgsm; tang2023large).
+Other studies find that models favor surface-level reasoning patterns over logical validity.
+Moreover, reasoning performance scales poorly with task difficulty, with models over-elaborating on simple problems and failing on harder ones (shojaee2025illusion).
+Concerns about faithfulness further arise from intervention-based analyses showing that final answers often remain unchanged when intermediate steps are corrupted or removed (lanham2023measuring), an effect referred to as the illusion of transparency (chen2025reasoning; bentham2024chainofthought).
+The opposing perspectives on CoT reasoning call for a systematic understanding of why and when CoT reasoning succeeds or fails.
 2.3 OOD Generalization of LLMs
-Out-of-distribution (OOD) generalization, where test inputs differ from training data, remains a key challenge in machine learning, particularly for large language models (LLMs)(Yang et al., 2024, 2023; Budnikov et al., 2025; Zhang et al., 2024b).
-Recent studies show that LLMs prompted to learn novel functions often revert to similar functions encountered during pretraining (Wang et al., 2024; Garg et al., 2022).
-Likewise, LLM generalization frequently depends on mapping new problems onto familiar compositional structures (Song et al., 2025).
-CoT prompting improves OOD generalization (Wei et al., 2022), with early work demonstrating length generalization for multi-step problems beyond training distributions (Yao et al., 2025; Shen et al., 2025).
-However, this ability is not inherent to CoT and heavily depends on model architecture and training setups.
-For instance, strong generalization in arithmetic tasks was achieved only when algorithmic structures were encoded into positional encodings (Cho et al., 2024).
-Similarly, finer-grained CoT demonstrations during training boost OOD performance, highlighting the importance of data granularity (Wang et al., 2025a).
-Theoretical and empirical evidence shows that CoT generalizes well only when test inputs share latent structures with training data; otherwise, performance declines sharply (Wang et al., 2025b; Li et al., 2025a).
-Despite its promise, CoT still struggles with genuinely novel tasks or formats.
-In the light of these brilliant findings, we propose rethinking CoT reasoning through a data distribution lens: decomposing CoT into task, length, and format generalization, and systematically investigating each in a controlled setting.
-Figure 2: Framework of DataAlchemy. It creates an isolated and controlled environment to train LLMs from scratch and probe the task, length, and format generalization.
-3 The Data Distribution Lens
-We propose a fundamental reframing to understand what CoT actually represents.
-We hypothesize that the underlying mechanism is better understood through the lens of data distribution: rather than executing explicit reasoning procedures, CoT operates as a pattern-matching process that interpolates and extrapolates from the statistical regularities present in its training distribution.
-Specifically, we posit that CoT’s success stems not from a model’s inherent reasoning capacity, but from its ability to generalize conditionally to out-of-distribution (OOD) test cases that are structurally similar to in-distribution exemplars.
-To formalize this view, we model CoT prompting as a conditional generation process constrained by the distributional properties of the training data.
-Let $\mathcal{D}_{\text{train}}$ denote the training distribution over input-output pairs $(x,y)$, where $x$ represents a reasoning problem and $y$ denotes the solution sequence (including intermediate reasoning steps).
-The model learns an approximation $f_{\theta}(x)\approx y$ by minimizing empirical risk over samples drawn from $\mathcal{D}_{\text{train}}$.
-Let the expected training risk be defined as:
-$R_{\text{train}}(f_{\theta})=\mathbb{E}_{(x,y)\sim\mathcal{D}_{\text{train}}}[\ell(f_{\theta}(x),y)],$ (1)
-where $\ell$ is a task-specific loss function (e.g., cross-entropy, token-level accuracy).
-At inference time, given a test input $a_{\text{test}}$ sampled from a potentially different distribution $\mathcal{D}_{\text{test}}$, the model generates a response $y_{\text{test}}$ conditioned on patterns learned from $\mathcal{D}_{\text{train}}$.
-The corresponding expected test risk is:
-$R_{\text{test}}(f_{\theta})=\mathbb{E}_{(x,y)\sim\mathcal{D}_{\text{test}}}[\ell(f_{\theta}(x),y)].$ (2)
-The degree to which the model generalizes from $\mathcal{D}_{\text{train}}$ to $\mathcal{D}_{\text{test}}$ is governed by the distributional discrepancy between the two, which we quantify using divergence measures:
-Definition 3.1 (Distributional Discrepancy).
-Given training distribution $\mathcal{D}_{\text{train}}$ and test distribution $\mathcal{D}_{\text{test}}$, the distributional discrepancy is defined as:
-$\Delta(\mathcal{D}_{\text{train}},\mathcal{D}_{\text{test}})=\mathcal{H}(\mathcal{D}_{\text{train}}\parallel\mathcal{D}_{\text{test}})$ (3)
-where $\mathcal{H}(\cdot\parallel\cdot)$ is a divergence measure (e.g., KL divergence, Wasserstein distance) that quantifies the statistical distance between the two distributions.
-Theorem 3.1 (CoT Generalization Bound).
-Let $f_{\theta}$ denote a model trained on $\mathcal{D}_{\text{train}}$ with expected training risk $R_{\text{train}}(f_{\theta})$.
-For a test distribution $\mathcal{D}_{\text{test}}$, the expected test risk $R_{\text{test}}(f_{\theta})$ is bounded by:
-$R_{\text{test}}(f_{\theta})\leq R_{\text{train}}(f_{\theta})+\Lambda\cdot\Delta(\mathcal{D}_{\text{train}},\mathcal{D}_{\text{test}})+\mathcal{O}\left(\sqrt{\frac{\log(1/\delta)}{n}}\right)$ (4)
-where $\Lambda>0$ is a Lipschitz constant that depends on the model architecture and task complexity, $n$ is the training sample size, and the bound holds with probability $1-\delta$, where $\delta$ is the failure probability.
-The proof is provided in Appendix A.1
-Building on this data distribution perspective, we identify three critical dimensions along which distributional shifts can occur, each revealing different aspects of CoT’s pattern-matching nature: ➊ Task generalization examines how well CoT transfers across different types of reasoning tasks.
-Novel tasks may have unique elements and underlying logical structure, which introduces distributional shifts that challenge the model’s ability to apply learned reasoning patterns.
-➋ Length generalization investigates CoT’s robustness to reasoning chains of varying lengths.
-Since training data typically contains reasoning sequences within a certain length range, test cases requiring substantially longer or shorter reasoning chains represent a form of distributional shift along the sequence length dimension.
-This length discrepancy could result from the reasoning step or the text-dependent solution space.
-➌ Format generalization explores how sensitive CoT is to variations in prompt formulation and structure.
-Due to various reasons (e.g., sophistical training data or diverse background of users), it is challenging for LLM practitioners to design a golden prompt to elicit knowledge suitable for the current case.
-Their detailed definition and implementation are given in subsequent sections.
-Each dimension provides a unique lens for understanding the boundaries of CoT’s effectiveness and the mechanisms underlying its apparent reasoning capabilities.
-By systematically varying these dimensions in controlled experimental settings, we can empirically validate our hypothesis that CoT performance degrades predictably as distributional discrepancy increases, thereby revealing its fundamental nature as a pattern-matching rather than reasoning system.
-4 DataAlchemy: An Isolated and Controlled Environment
-To systematically investigate the influence of distributional shifts on CoT reasoning capabilities, we introduce DataAlchemy, a synthetic dataset framework designed for controlled experimentation.
-This environment enables us to train language models from scratch under precisely defined conditions, allowing for rigorous analysis of CoT behavior across different OOD scenarios.
-The overview is shown in Figure 2.
+Out-of-distribution (OOD) generalization remains a central challenge in machine learning (yang2024generalized; yang2023out; budnikov2025generalization; zhang2024can).
+Prior work shows that pre-trained models face challenges in adapting to new settings when prompted to learn novel functions (wang2024can; garg2022can).
+Researchers reveal CoT prompting can partially improve OOD generalization ability, especially for tasks that require long reasoning (yao2025unveiling; shen2025codi).
+However, other work claims such gains are not intrinsic.
+For example, strong arithmetic generalization emerges only when algorithmic biases are encoded in positional representations (cho2024position), and finer-grained CoT supervision during training substantially improves OOD performance (wang2025chain).
+Recent studies further indicate that LLM generalizes reliably when common latent structures are shared across distributions (wang2025theoretical; li2025training).
+In light of these brilliant findings, we propose rethinking CoT reasoning through a data distribution lens, dissecting CoT reasoning into task, length, and format, and systematically investigating through controlled experiments.
+We further provide a Comparison with representative work in Appendix A.4.
+3 The Proposed Data Distribution Lens
+We propose the data distribution lens to understand why and when CoT reasoning succeeds or fails.
+We hypothesize that CoT reasoning reflects a structured inductive bias learned from in-distribution data, enabling models to conditionally generate reasoning trajectories that approximate those observed during training.
+As such, the effectiveness of CoT reasoning is fundamentally governed by the nature and degree of distribution discrepancy between training data and test queries (rather than by model architecture or scale).
+To formalize this view, we first introduce notation for the training and test distributions.
+Let $\mathcal{D}_{\text{train}}$ denote the training distribution over input-output pairs $(x,y)$, where $x$ represents a reasoning problem and $y$ denotes the solution sequence (including intermediate reasoning traces).
+During training, the model learns a parametric mapping $f_{\theta}(x)\approx y$ by minimizing the empirical training risk
+$\hat{R}_{\text{train}}(f_{\theta})=\frac{1}{n}\sum_{i=1}^{n}\ell\bigl(f_{\theta}(x_{i}),y_{i}\bigr)$ (1)
+where $(x_{i},y_{i})\sim\mathcal{D}_{\text{train}}$ are i.i.d. samples and $\ell$ is a loss function (e.g., cross-entropy). The corresponding expected (population) training risk is
+$R_{\text{train}}(f_{\theta})=\mathbb{E}_{(x,y)\sim\mathcal{D}_{\text{train}}}\bigl[\ell(f_{\theta}(x),y)\bigr]$ (2)
+At inference time, given a test query sampled from a potentially different distribution $\mathcal{D}_{\text{test}}$, the model generates a response. The expected test risk is
+$R_{\text{test}}(f_{\theta})=\mathbb{E}_{(x,y)\sim\mathcal{D}_{\text{test}}}\bigl[\ell(f_{\theta}(x),y)\bigr]$ (3)
+Definition 3.1 (Distribution Discrepancy).
+Given training distribution $\mathcal{D}_{\text{train}}$ and test distribution $\mathcal{D}_{\text{test}}$, we define the distribution discrepancy as
+$\Delta(\mathcal{D}_{\text{train}},\mathcal{D}_{\text{test}}):=\operatorname{TV}(\mathcal{D}_{\text{train}}\mathcal{D}_{\text{test}})$ (4)
+where $\operatorname{TV}(P,Q)$ is the total variation distance,
+$\operatorname{TV}(P,Q):=\sup_{A}|P(A)-Q(A)|=\frac{1}{2}\int|dP-dQ|$ (5)
+Theorem 3.1 (Generalization Bound).
+Assume the loss is bounded, i.e., for all $(x,y)$, $0\leq\ell(f_{\theta}(x),y)\leq B$.
+Let $\{(x_{i},y_{i})\}_{i=1}^{n}$ be i.i.d.
+samples from $\mathcal{D}_{\text{train}}$ and let $\hat{R}_{\text{train}}(f_{\theta})$ be the empirical training risk defined above.
+Then for any $\delta\in(0,1)$, with probability at least $1-\delta$ over the draw of the training sample, the expected test risk satisfies
+$R_{\text{test}}(f_{\theta})\!\leq\!\hat{R}_{\text{train}}(f_{\theta})\!+\!2B\,\Delta(\mathcal{D}_{\text{train}},\mathcal{D}_{\text{test}})\!+\!B\sqrt{\frac{\log(1/\delta)}{2n}}.$ (6)
+The proof is provided in Appendix H.1.
+Theorem 3.1 provided a theoretical foundation for the data distribution lens. Guided by it, we identify three critical dimensions along which distribution shifts can occur: task, length, and format.
+$\Delta(\mathcal{D}_{\text{train}},\mathcal{D}_{\text{test}})=\Phi\left(\Delta_{\text{task}},\Delta_{\text{length}},\Delta_{\text{format}}\right)$ (7)
+where $\Phi$ is a monotonically increasing composition function that aggregates all discrepancies.
+$\Delta_{\text{task}}$, $\Delta_{\text{length}}$, and $\Delta_{\text{format}}$ measure the distribution discrepancy introduced by unseen tasks, various lengths, and prompt format variants.
+Figure 2: Framework of DataAlchemy.
+DataAlchemy provides an abstract representation system that distills various real-world NLP tasks into key components: atoms, elements, and transformations.
+By varying these components, we curate data that exhibits various distribution discrepancies following task, length, and format generalization.
+DataAlchemy achieves full and fine-grained control over the entire evaluation pipeline.
+Later, we train models from scratch to avoid data leakage and employ controlled experiments to rigorously test the hypotheses.
+4 DataAlchemy: A Controllable Environment
+To empirically validate the data distribution lens, we introduce DataAlchemy, an abstract, controllable, and clean environment.
+It distills real-world NLP tasks into basic atoms, elements, and transformations as illustrated in Figure 2.
 4.1 Basic Atoms and Elements
-Let $\mathcal{A}=\{\texttt{A},\texttt{B},\texttt{C},\ldots,\texttt{Z}\}$ denote the alphabet of 26 basic atoms. An element $\mathbf{e}$ is defined as an ordered sequence of atoms:
-$\mathbf{e}=(a_{0},a_{1},\ldots,a_{l-1})\quad\text{where}\quad a_{i}\in\mathcal{A},\quad l\in\mathbb{Z}^{+}$ (5)
-This design provides a versatile manipulation for the size of the dataset $\mathcal{D}$ (i.e., $|\mathcal{D}|=|\mathcal{A}|^{l}$) by varying element length $l$ to train language models with various capacities.
-Meanwhile, it also allows us to systematically probe text length generalization capabilities.
+We abstract tokens in the real-world NLP tasks into basic atoms represented by an alphabet of 26 $\mathcal{A}=\{\texttt{A},\texttt{B},\texttt{C},\ldots,\texttt{Z}\}$.
+Based on atoms, we further construct element $\mathbf{e}$ as an ordered sequence of atoms with length $l$, reflecting the text space (considering the text consists of tokens):
+$\mathbf{e}=(a_{0},a_{1},\ldots,a_{l-1})\quad\text{where}\ a_{i}\in\mathcal{A},\ l\in\mathbb{Z}^{+}$ (8)
+Note that we can construct at most $|\mathcal{A}|^{l}$ distinct elements, which provides a versatile approach for data curation by manipulating element length $l$.
 4.2 Transformations
-A transformation is an operation that operates on elements $F:\mathbf{e}\rightarrow\hat{\mathbf{e}}$.
-In this work, we consider two fundamental transformations: the ROT Transformation and the Cyclic Position Shift.
+Similarly, we abstract operations LLM performed on text in the real world (e.g., summarize, paraphrase, and reasoning) as transformations that operate on elements $F:\mathbf{e}\rightarrow\hat{\mathbf{e}}$.
+In this work, we mainly instantiate two fundamental transformations: the ROT Transformation and the Cyclic Position Shift.
+Additional transformations are considered in the Appendix F.1 to avoid bias.
 To formally define the transformations, we introduce a bijective mapping $\phi:\mathcal{A}\to\mathbb{Z}_{26}$, where $\mathbb{Z}_{26}=\{0,1,\ldots,25\}$, such that $\phi(c)$ maps a character to its zero-based alphabetical index.
@@
 The compositional transformation $f_{\text{S}}$ for the sequence $S$ is the function composition:
-$f_{\text{S}}=f_{k}\circ f_{k}\circ\cdots\circ f_{1}$ (8)
+$f_{\text{S}}=f_{1}\circ f_{2}\circ\cdots\circ f_{k}$ (11)
 The resulting element $\hat{\mathbf{e}}$ is obtained by applying the operations sequentially to an initial element $\mathbf{e}$:
 $\hat{\mathbf{e}}=f_{k}(f_{k-1}(\ldots(f_{1}(\mathbf{e}))\ldots))$ (12)
-This design enables the construction of arbitrarily complex transformation chains by varying the type, parameters, order, and length of operations within the sequence.
-At the sample time, we can naturally acquire the COT reasoning step by decomposing the intermediate process:
-$\underbrace{f_{\text{S}}(\mathbf{e}):}_{\text{Query}}\quad\underbrace{\mathbf{e}\xrightarrow{f_{1}}\mathbf{e}^{(1)}\xrightarrow{f_{2}}\mathbf{e}^{(2)}\cdots\xrightarrow{f_{k-1}}\mathbf{e}^{(k-1)}\xrightarrow{f_{k}}}_{\text{COT reasoning steps}}\underbrace{\boxed{\hat{\mathbf{e}}}}_{\text{Answer}}$ (10)
+This design enables the construction of arbitrary transformations with the type, parameters, order, and length.
+At the same time, we can naturally acquire the CoT reasoning step by decomposing the intermediate process:
+$\underbrace{f_{\text{S}}(\mathbf{e}):}_{\text{Query}}\quad\underbrace{\mathbf{e}\xrightarrow{f_{1}}\mathbf{e}^{(1)}\xrightarrow{f_{2}}\mathbf{e}^{(2)}\cdots\xrightarrow{f_{k-1}}\mathbf{e}^{(k-1)}\xrightarrow{f_{k}}}_{\text{Reasoning traces}}\underbrace{\boxed{\hat{\mathbf{e}}}}_{\text{Answer}}$ (13)
+Illustrative examples of atoms, elements, and transformations are detailed in Appendix B.
 4.3 Environment Setting
-Through systematic manipulation of elements and transformations, DataAlchemyoffers a flexible and controllable framework for training LLMs from scratch, facilitating rigorous investigation of diverse OOD scenarios.
-Without specification, we employ a decoder-only language model GPT-2 (Radford et al., 2019) with a configuration of 4 layers, 32 hidden dimensions, and 4 attention heads.
-We utilize a Byte-Pair Encoding (BPE) tokenizer.
-Both LLMs and the tokenizer follow the general modern LLM pipeline.
+Through systematic manipulation of elements and transformations, DataAlchemy, we can train and probe various LLMs under various tasks, lengths, and format distributions.
+In the controlled experiment, we employ decoder-only LLMs with GPT and LLaMA architectures and parameter sizes ranging from 62K to 3B when training from scratch.
+In the real-world experiments, we utilize two state-of-the-art (SOTA) LLMs: LLaMA3-8B-Instruct dubey2024LLaMA and Qwen3-14B yang2025qwen3.
+We construct elements with 2 to 6 basic atoms, which produce 676 to 308,915,776 data samples.
+We initialize the two transformations $f_{1}=f_{\text{rot}}(e,13)$ and $f_{2}=f_{\text{pos}}(e,1)$.
+We consider both hard metrics, i.e., exact match rate, and soft metrics, i.e., Levenshtein distance (edit distance) (yujian2007normalized), and BLEU score (papineni2002bleu) for evaluation.
+To enable a fine-grained analysis, we evaluate reasoning traces, the final answer, and the full chain in the LLM response.
+Detailed environment setting and implementation are provided in Appendix D.
+Figure 3: Transformation generalization under different distribution discrepancies. The efficacy of CoT reasoning decreases as task distribution discrepancy increases.
+5 Task Generalization
+To investigate the extent to which CoT reasoning can handle tasks under various distribution discrepancies, we design task generalization experiments.
+As we discussed in Section 4, we decompose tasks into a combination of various transformations and elements.
+Therefore, we consider task generalization from two dimensions: transformation generalization and element generalization.
+5.1 Transformation Generalization
+Experiment setup.
+To formulate different distribution discrepancies for task generalization, we design the following progressive scenarios based on the proposed measurement (detailed in Appendix E.1).
+(i) In-Distribution (ID).
+The transformations in the test set are identical to those observed during training, e.g., $f_{1}\circ f_{1}\rightarrow f_{1}\circ f_{1}$.
+(ii) Composition (CMP).
+Test samples comprise novel compositions, where basic transformations are observed during training, e.g., ${f_{1}\circ f_{1},f_{1}\circ f_{2},f_{2}\circ f_{1}}\rightarrow f_{2}\circ f_{2}$.
+(iii) Partial Out-of-Distribution (POOD): Test queries include compositions involving both seen and unseen basic transformations, e.g., $f_{1}\circ f_{1}\rightarrow f_{1}\circ f_{2}$.
+(iv) Out-of-Distribution (OOD).
+The test set contains entirely novel transformations (compositions) in training, e.g., $f_{1}\circ f_{1}\rightarrow f_{2}\circ f_{2}$.
+The illustrative examples for transformation generalization under different scenarios are provided in Appendix C.1.1.
+Table 1: Full chain evaluation under different scenarios on transformation generalization.
+Scenarios Exact Match (%) Edit Distance BLEU Score
+ID 100.00 0 1
+CMP 0.01 0.1326 0.6867
+POOD 0.00 0.1671 0.4538
+OOD 0.00 0.2997 0.2947
+Findings.
+Figure 3 illustrates the performance of the full chain under different distribution discrepancies.
+We can observe that, in general, the effectiveness of CoT reasoning decreases as the distribution discrepancy increases, which directly validates the data distribution lens.
+As shown in Table 1, CoT reasoning achieves satisfactory performance in the ID (exact match: 100%) scenario, while it degrades in CMP (0.01%), POOD (0%), and OOD (0%) scenarios.
+Diving into fine-grained analysis, as demonstrated in Table 2, we find that the success of CoT reasoning is attributed to the replicating pattern in the training data, as indicated by the inconsistency in reasoning and answers.
+For instance, when an unseen transformation $f_{1}\circ f_{1}$ is present, LLMs attempt to generalize based on the most similar transformation (i.e., $f_{1}\circ f_{2}$) seen during training, which leads to correct reasoning paths yet incorrect answers.
+As the commutativity of the transforms, generalization from $f_{1}\circ f_{2}$ to $f_{2}\circ f_{1}$ or vice versa allows LLMs to produce incorrect paths yet correct answers, which reflect the unfaithfulness and pattern-matching nature of CoT reasoning.
+Additional analysis and illustrative examples are provided in Appendix F.1 and G.1.
+Table 2: Fine-grained analysis for CoT reasoning on transformation generalization based on exact match.
+Transformation (Train $\rightarrow$ Test) Reasoning Answer Full Chain
+$\{f_{1}\circ f_{1},f_{1}\circ f_{2},f_{2}\circ f_{1}\}\rightarrow f_{2}\circ f_{2}$ 100.00 0.01 0.01
+$\{f_{1}\circ f_{2},f_{2}\circ f_{1},f_{2}\circ f_{2}\}\rightarrow f_{1}\circ f_{1}$ 100.00 0.01 0.01
+$f_{1}\circ f_{2}\rightarrow f_{2}\circ f_{1}$ 0.00 100.00 0.00
+$f_{2}\circ f_{1}\rightarrow f_{1}\circ f_{2}$ 0.00 100.00 0.00
+Experiment setup. To further probe when CoT reasoning can adapt to unseen transformations, we conduct supervised fine-tuning (SFT) experiments to incorporate a portion $\lambda$ of unseen data.
+Findings.
+As shown in Figure 4, we can find that generally a very small portion ($\lambda=1.5e{-4}$) of data can make the model quickly generalize to unseen transformations.
+The less discrepancy between the training and testing data, the easier the model can generalize, highlighting the role of similar patterns that appear in the training data.
+5.2 Element Generalization
+Figure 4: Effectiveness of SFT. A small portion of unseen data helps CoT reasoning to quickly generalize.
+Following a pipeline similar to transformation generalization, we investigate how CoT reasoning handles elements under various distribution discrepancies.
+Findings observed also support the proposed data distribution.
+The detailed experiment design and analysis can be found in Appendix F.2.
+6 Length Generalization
+To study how CoT reasoning can operate on varying lengths, we design a length generalization experiment.
+Following the same intuition as task generalization, we also formulate length generalization as two perspectives: text length (i.e., element length) generalization and reasoning step (i.e., transformation length) generalization.
+6.1 Text Length Generalization
+Experiment setup.
+The text length distribution discrepancy can be measured by element length difference, detailed in Appendix E.2.
+We train LLMs on the dataset with text length $l=4$ while fixing other factors and evaluate the performance on a variety of lengths (e.g., from $l=2$ to $l=6$).
+We provide illustrative examples for text length generalization in Appendix C.2.1.
+Figure 5: Text length generalization under distribution discrepancies. Increasing distribution shifts in the text length lead to degraded CoT reasoning performance.
+Findings.
+As illustrated in the Figure 5, CoT reasoning produces excellent results under in-distribution scenarios ($l=4$), while its performance degrades as discrepancies in the text length distribution increase, which serves as evidence for the data distribution lens.
+When we further analyze the exact match in the Table 5, the CoT reasoning failed to directly generate test cases for those lengths, even with a mild distribution shift (e.g., $l=3$ or $l=5$).
+Examples in Appendix G.2.2 indicate that LLMs attempt to produce CoT reasoning with the same length as the training data by adding or removing tokens when processing unseen text length.
+We further consider the effect of different padding strategies in Appendix F.3.
+6.2 Reasoning Step Generalization
+Experiment setup.
+Reasoning steps are determined by the number of basic transformations $k$ in the compositional transformation.
+We mix the data with various reasoning steps (e.g., $k=1,2,3$).
+By adjusting the mix ratio while maintaining the data size, we create different distribution discrepancies.
+Examples of reasoning step generalization are detailed in Appendix C.2.2.
+Figure 6: Reasoning step generalization under distribution discrepancies. Performance of CoT reasoning systematically varies with training data components.
+Findings.
+As showcased in Figure 6, when we adjust the component of training data, the performance of CoT changed accordingly.
+For instance, increasing the ratio of $k=1$ data will enhance performance on one-step reasoning while compromising two-step reasoning, which supports our hypothesis.
+Notably, considering extreme cases where the mix ratio is 0 or 1.0, the CoT reasoning achieves good performance in the covered reasoning step but fails to generalize to unseen cases, indicating its fragility when encountering distribution shifts.
+7 Format Generalization
+To research the robustness of CoT reasoning when surface-level variations appear in test queries, we designed the format generalization.
+Figure 7: Format generalization under distribution discrepancies. Testing performance degrades with various noise levels and in different applied areas.
+Experiment setup.
+To introduce the distribution discrepancy at a formal level, we proposed distribution measurement (detailed in Appendix E.3) and consider four distinct perturbation modes to simulate a scenario in the real world.
+(i) Insert.
+One noise token is inserted; (ii) Delete.
+One original token is deleted; (iii) Modify: One original token is replaced with a noise token; and (iv) Hybrid: it combines the above-mentioned perturbation methods.
+We apply four perturbations with the noise level of $p$ on different areas (e.g., elements, transformations, and prompts) of test queries.
+We further elaborate on the format generalization using illustrative examples in Appendix C.3.
+Findings.
+As observed in Figure 7, introducing perturbation will compromise the effectiveness of CoT reasoning, and the degree depends on the noise level (i.e., distributional shift), which echoes the data distribution lens.
+Among different perturbation methods, insertion makes the greatest difference.
+Considering different areas applied, the elements and transformation play an important role, whereas the changes to other tokens have a lesser effect on the results, which aligns with intuition.
+8 Generality of Data Distribution Lens
+To probe the generality of the data distribution length, we design experiments using LLMs with various architectures, sizes, and temperatures.
+8.1 Internal Validity
+Experiment setup. For rigor, we conduct the experiments of task, length, and format generalization by training LLMs with GPT and LLaMA architectures and sizes ranging from 62K to 3B.
+Findings.
+As illustrated in Figure 8, CoT reasoning produced by LLM with different sizes and architectures behaves similarly when encountering distribution shifts on task, length, and format generalization, highlighting the good internal validity.
+We further study the effect of temperature and the role of SFT with different model sizes in Appendix F.4.
+Figure 8: Task, length, and format generalization of LLMs with different settings.
+The data distribution lens is invariant across LLMs with various sizes and architectures.
+Results under more settings are provided in Figure 15.
+8.2 External Validity
+The key to the external validity of the data distribution lens is to identify the distribution discrepancy between training data and test queries, which makes direct evaluation extremely challenging due to the opacity of the training data used by SOTA LLMs.
+However, this problem can be alleviated if we can curate data unseen during training and then use it to fine-tune LLMs.
+By interfering with data generated by DataAlchemy, where LLMs produce totally random answers, we confirm the vanity of the proposed pipelines.
+Experiment setup. We conduct the experiments of task, length, and format generalization fine-tuning two SOTA LLMs: LLaMA3-8B-Instruct dubey2024LLaMA and Qwen3-14B yang2025qwen3.
+Findings.
+As shown in Figure 9, the performance of SOTA LLMs exhibits similar trends to pre-trained models in DataAlchemy across task, length, and format generalization, indicating the external validity of the distribution lens.
+Additional results and analysis are provided in Appendix F.6.
+Figure 9: Task, length, and format generalization of SOTA LLMs. The data-distribution lens is valid.
+9 Discussion and Implication
+Through rigorous experiments, we demonstrate that CoT reasoning is effective when applied to (near) in-distribution data, but becomes fragile and prone to failure even under moderate distribution shifts.
+What appears to be structured reasoning can be a mirage, emerging from memorized or interpolated patterns in the training data rather than logical inference.
+Our work carries important implications for both LLM researchers and practitioners, which are further discussed in Appendix I.
+10 Conclusion
+We examine the CoT reasoning of LLMs through the data distribution lens, revealing that the perceived structured reasoning capability largely arises from inductive biases shaped by in-distribution training data, whose effectiveness is bounded by distribution discrepancies.
+We propose a fully controllable framework, DataAlchemy, and systematically probe CoT reasoning with distribution discrepancies introduced by tasks.
+length and format.
+Comprehensive experiments confirm that the data distribution is invariant across LLMs with different architectures and sizes.
+We hope DataAlchemy can serve as a platform where researchers can rigorously explore the nature of LLMs, inspiring the discovery of universal principles.
+Limitations
+While our work offers a rigorous, controlled investigation into the nature of Chain-of-Thought (CoT) reasoning, we acknowledge several limitations that provide avenues for future research:
+(i) Synthetic environment vs.
+natural language complexity.
+Our controlled experiments on the abstract environment DataAlchemy, which distills real-world language tasks into symbolic atoms, elements, and transformations.
+While this abstraction enables full and fine-grained control over distribution factors and avoids data leakage, it may inevitably not fully capture the semantic richness, ambiguity, and compositional diversity present in natural language.
+While external validity of the proposed data distribution is confirmed by real-world SOTA LLMs, the observed brittleness of CoT reasoning under distribution shifts may manifest more stealthily, sophistically, and task-dependently in more complex real-world settings.
+(ii) Distribution discrepancy measurement and data opacity.
+Although we evaluate a wide range of model architectures, sizes, and temperatures, including both models trained from scratch and state-of-the-art pretrained LLMs, the training data distributions of commercial or large proprietary models remain uncovered due to the opaque nature of training data and model weights.
+As a result, fully estimating the distribution discrepancy between pretraining data and test queries is inherently challenging, limiting the precision with which our data distribution lens can be quantitatively validated in fully realistic and transparent scenarios.
+(iii) Scope of generalization dimensions.
+We focused our analysis on three specific dimensions of generalization: task, length, and format.
+While these cover a broad spectrum of OOD scenarios, we did not explicitly model other forms of distribution shift, such as cross-lingual transfer, multi-modal reasoning, or shifts in cultural context.
+Ethical Considerations
+This work studies the reasoning behavior of large language models and does not involve human subjects, personal data, or user-generated content.
+All experiments are conducted on synthetic or publicly available benchmarks and models, and models trained from scratch use data generated entirely within the proposed framework, avoiding issues of privacy, consent, or data misuse.
+Our findings highlight that CoT reasoning can produce fluent yet logically inconsistent or unfaithful reasoning traces when models are evaluated outside their training distributions.
+This has ethical implications for the deployment of LLMs in high-stakes applications such as education, healthcare, law, and scientific decision-making, where users may over-trust seemingly coherent reasoning explanations.
+We emphasize that the presence of a detailed reasoning trace should not be equated with correctness, reliability, or genuine understanding.
+By exposing the limitations and fragility of CoT reasoning, this work aims to promote more responsible use of LLMs and encourage the research community to develop evaluation protocols and modeling approaches that better reflect true generalization and reasoning capabilities.
+We believe that transparency about these limitations is essential for preventing misuse and misinterpretation of LLM-generated reasoning.
+[…]
+Appendix A Extended Related Work and Comparison
+A.1 LLM Prompting and CoT
+Chain-of-Thought (CoT) prompting revolutionized how we elicit reasoning from Large Language Models by decomposing complex problems into intermediate steps [wei2022chain].
+By augmenting few-shot exemplars with reasoning chains, CoT showed substantial performance gains on various tasks [xu2024faithful, imani2023mathprompter, wei2022chain].
+Building on this, several variants emerged.
+Zero-shot CoT triggers reasoning without exemplars using instructional prompts [kojima2022large], and self-consistency enhances performance via majority voting over sampled chains [wang2023selfconsistency].
+To reduce manual effort, Auto-CoT generates CoT exemplars using the models themselves [zhang2023automatic].
+Beyond linear chains, Tree-of-Thought (ToT) frames CoT as a tree search over partial reasoning paths [yao2023tree], enabling lookahead and backtracking.
+SymbCoT combines symbolic reasoning with CoT by converting problems into formal representations [xu2024faithful].
+Recent work increasingly integrates CoT into the LLM inference process, generating long-form CoTs [jaech2024openai, team2024qwq, guo2025deepseek, team2025kimi].
+This enables flexible strategies like mistake correction, step decomposition, reflection, and alternative reasoning paths [yeo2025demystifying, chen2025towards].
+The success of prompting techniques and long-form CoTs has led many to view them as evidence of emergent, human-like reasoning in LLMs.
+In this work, we investigate whether CoT reflects genuine reasoning or merely pattern interpolation.
+A.2 Discussion on Illusion of LLM Reasoning
+While Chain-of-Thought prompting has led to impressive gains on complex reasoning tasks, a growing body of work has started questioning the nature of these gains [stechly2024chain].
+One major line of research highlights the fragility of CoT reasoning.
+Minor and semantically irrelevant perturbations such as distractor phrases or altered symbolic forms can cause significant performance drops in state-of-the-art models [mirzadehgsm, tang2023large].
+Models often incorporate such irrelevant details into their reasoning, revealing a lack of sensitivity to salient information.
+Other studies show that models prioritize the surface form of reasoning over logical soundness; in some cases, longer but flawed reasoning paths yield better final answers than shorter, correct ones [bentham2024chainofthought].
+Similarly, performance does not scale with problem complexity as expected—models may overthink easy problems and give up on harder ones [shojaee2025illusion].
+Another critical concern is the faithfulness of the reasoning process.
+Intervention-based studies reveal that final answers often remain unchanged even when intermediate steps are falsified or omitted [lanham2023measuring], a phenomenon dubbed the illusion of transparency [bentham2024chainofthought, chen2025reasoning].
+Together, these findings suggest that LLMs are not principled reasoners but rather sophisticated simulators of reasoning-like text.
+However, a systematic understanding of why and when CoT reasoning succeeds or fails is still a mystery.
+A.3 OOD Generalization of LLMs
+Out-of-distribution (OOD) generalization, where test inputs differ from training data, remains a key challenge in machine learning, particularly for large language models (LLMs)[yang2024generalized, yang2023out, budnikov2025generalization, zhang2024can].
+Recent studies show that LLMs prompted to learn novel functions often revert to similar functions encountered during pretraining [wang2024can, garg2022can].
+Likewise, LLM generalization frequently depends on mapping new problems onto familiar compositional structures [song2025out].
+CoT prompting improves OOD generalization [wei2022chain], with early work demonstrating length generalization for multi-step problems beyond training distributions [yao2025unveiling, shen2025codi].
+However, this ability is not inherent to CoT and heavily depends on model architecture and training setups.
+For instance, strong generalization in arithmetic tasks was achieved only when algorithmic structures were encoded into positional encodings [cho2024position].
+Similarly, finer-grained CoT demonstrations during training boost OOD performance, highlighting the importance of data granularity [wang2025chain].
+Theoretical and empirical evidence shows that CoT generalizes well only when test inputs share latent structures with training data; otherwise, performance declines sharply [wang2025theoretical, li2025training].
+Despite its promise, CoT still struggles with genuinely novel tasks or formats.
+In the light of these brilliant findings, we propose rethinking CoT reasoning through a data distribution lens: decomposing CoT into task, length, and format generalization, and systematically investigating through controlled experiments.
+Figure 10: Comparison with representative evaluation method.
+DataAlchemy distills real-world NLP problems, allowing training LLMs from scratch to avoid data leakage issues and study CoT reasoning through rigorous controlled experiments.
+A.4 Comparison with Representative Work
+Recent research attempts to create a fine-grained evaluation by examining both the CoT reasoning trace and the final answer through controlled experiments.
+However, they suffer from: (i) Narrowly defined settings: focusing on specific tasks, domains, or LLMs, thereby overlooking the common characteristics across tasks and LLMs.
+(ii) Data entanglement: most evaluations are conducted on real-world tasks and models, where the complexity precludes fully controlled experiments.
+(iii) Data leakage: LLMs train makes the best of all the data, including benchmarks, undermining the effectiveness and validity of evaluations, which is illustrated in Figure 10.
+To scientifically and rigorously examine CoT reasoning, a new evaluation framework is required.
+An ideal framework should satisfy the following criteria: (i) Abstract representation: it should abstract and unify diverse NLP tasks and LLMs while retaining their essential properties.
+(ii) Fully controlled experiment: it should enable a full and fine-grained control of both tasks (e.g., complexity) and LLMs (e.g., size and architecture), enabling rigorous study of different factors through controlled experiments.
+(iii) Training from scratch: it should offer scalable structural data to train LLMs from scratch, mitigating the data leakage and providing clean evaluations.
+Additionally, our proposed framework enables the study of why and when CoT reasoning succeeds or fails.
+Appendix B Illustration of DataAlchemy Components
+B.1 Elements
+As defined in Section 4.1, an element $e$ is defined as an ordered sequence of atoms. The following is an element with 3 atoms:
+A P P
+The length of the elements can vary and the following is an element with 4 atoms:
+A P P L
+B.2 Transformations
+A transformation is defined as an operator acting on the elements, as detailed in Section 4.2. In this work, we employ three distinct transformations, denoted as $f_{1}$, $f_{2}$, and $f_{3}$.
+Specifically, $f_{1}$ represents an element-wise ROT-13 operation applied to the alphabet. The following example demonstrates the application of $f_{1}$ to the input APPL:
+A P P L [F1] <answer> N C C Y
+$f_{2}$ denotes a cyclic positional shift by one step. The following example illustrates the result of applying $f_{2}$ to the sequence APPL:
+A P P L [F2] <answer> P P L A
+$f_{3}$ denotes the sequence reversal operator. The following example displays the result of applying $f_{3}$ to APPL:
+A P P L [F3] <answer> L P P A
+B.3 Compositional Transformations
+Compositional transformations refer to imposing multiple transformations sequentially to an element. Below are the examples of transformations $f_{1}$ and $f_{2}$:
+A P P L [F1] [F2] <think> N C C Y [F2] <answer> C C Y N
+Appendix C Illustration of Generalization Tasks
+C.1 Illustration of Task Generalization
+As shown in Fig. 2, task is defined through element and transformation and hence the generalization of tasks can be decomposed to generalization on element and on transformation respectfully.
+C.1.1 Transformation Generalization
+We consider the transformation generalization on four aspects: in distribution (ID); compositional (CMP), partially out of distribution (POOD) and out of distribution (OOD).
+We use the element APPL to further demonstrate the four aspects in detail:
+In distribution
+refers to the scenario that the test transformations are identical to the training ones:
+Training:
+A P P L [F1] [F2] <think> N C C Y [F2] <answer> C C Y N
+Test:
+A P P L [F1] [F2] <think>
+Compositional
+refers to the scenario where the test transformations are the composition of the training ones
+Training:
+A P P L [F1] [F2] <think> N C C Y [F2] <answer> C C Y N
+A P P L [F2] [F1] <think> P P L A [F1] <answer> C C Y N
+A P P L [F2] [F2] <think> P P L A [F2] <answer> P L A P
+Test:
+A P P L [F1] [F1] <think>
+Partially out of distribution
+refers to the scenario where part of the compositional test transformations are seen during training while the entire compositional transformation is different from training
+Training:
+A P P L [F1] [F1] <think> N C C Y [F2] <answer> A P P L
+Test:
+A P P L [F1] [F2] <think>
+Out of distribution
+refers to the scenario where none of the compositional test transformations are seen during training.
+Training:
+A P P L [F2] [F2] <think> P P L A [F2] <answer> P L A P
+Test:
+A P P L [F1] [F1] <think>
+C.1.2 Element Generalization
+We consider the transformation generalization on three aspects: in distribution (ID); compositional (CMP) and out of distribution (OOD).
+We use the element APPL to further demonstrate the three aspects in detail:
+In distribution
+refers to the scenario that the test elements are identical to the training ones:
+Training:
+A P P L [F1] [F2] <think> N C C Y [F2] <answer> C C Y N
+Test:
+A P P L [F1] [F2] <think>
+Compositional
+refers to the scenario where the test elements have the same atoms as the training ones but with a different order
+Training:
+A P P L [F1] [F2] <think> N C C Y [F2] <answer> C C Y N
+Test:
+P A L P [F1] [F2] <think>
+Out of distribution
+refers to the scenario where the novel atoms show in the test elements
+Training:
+A P P L [F1] [F2] <think> N C C Y [F2] <answer> C C Y N
+Test:
+A P P Y [F1] [F2] <think>
+C.2 Illustration of Length Generalization
+Similar to above, we decompose the length generalization into text length generalization and reasoning step generalization.
+C.2.1 Illustration of Text Length Generalization
+Text length generalization involves different lengths of elements in the test set from those in the training set. Still, we use APPL as an example to illustrate text length generalization.
+Training:
+A P P L [F2] [F2] <think> P P L A [F2] <answer> P L A P
+Test:
+A P P L E [F1] [F2] <think>
+Test:
+A P P [F1] [F2] <think>
+C.2.2 Illustration of Reasoning Step Generalization
+Reasoning step generalization refers to the scenario where the number of transformations in the test compositional transformation is different from that in the training compositional transformation.
+Training:
+A P P L [F2] [F2] <think> P P L A [F2] <answer> P L A P
+Test:
+A P P L [F2] <think>
+Test:
+A P P L [F2] [F2] [F2] <think>
+C.3 Illustration of Format Generalization
+Format generalization refers to scenarios where the input formats are changed during testing.
+In this work, we consider three different mechanisms that the format gets changed: addition, deletion and modification: Insert.
+It refers to an additional unknown token that is added to the task.
+The following example shows how the APPL gets changed under addition:
+Training:
+A P P L [F2] [F2] <think> P P L A [F2] <answer> P L A P
+Test:
+A P <noise> P L [F2] [F2] <think>
+Delete refers to tokens that are deleted during the test. The following example shows how the APPL gets changed under delete:
+Training:
+A P P L [F2] [F2] <think> P P L A [F2] <answer> P L A P
+Test:
+A P L [F2] [F2] <think>
+Modify. It refers to tokens that are replaced by an unknown token during the test. The following example shows how the APPL gets changed under modification:
+Training:
+A P P L [F2] [F2] <think> P P L A [F2] <answer> P L A P
+Test:
+A <noise> P L [F2] [F2] <think>
+Appendix D Experiment Environment and Implementation Details
+D.1 Environment Setup
+We conduct controlled experiments using LLMs with different sizes (ranging from 62K to 3B) and different architectures (GPT and LLaMA) of LLMs; detailed hyperparameters are summarized in Table 3.
+For LLMs trained from scratch, we employ the AdamW optimiser in mixed precision (FP16).
+The default learning rate is $3\times 10^{-3}$, and the schedule follows a cosine decay with a 10% warm-up ratio.
+Training is conducted using a batch size of 1024, and each model is optimized for 10 epochs.
+A weight decay of 0.01 is applied, and gradient norms are clipped at 1.0.
 During the inference time, we set the temperature to 1e-5.
-For rigor, we also study LLMs with various parameters, architectures, and temperatures in Section 8.
-Details of the implementation are provided in the Appendix B.
-We consider that each element consists of 4 basic atoms, which produces 456,976 samples for each dataset with varied transformations and token amounts.
-We initialize the two transformations $f_{1}=f_{\text{rot}}(e,13)$ and $f_{2}=f_{\text{pos}}(e,1)$.
-We consider the exact match rate, Levenshtein distance (i.e., edit distance) (Yujian and Bo, 2007), and BLEU score (Papineni et al., 2002) as metrics and evaluate the produced reasoning step, answer, and full chain.
-Examples of the datasets and evaluations are shown in Appendix C
-5 Task Generalization
-Task generalization represents a fundamental challenge for CoT reasoning, as it directly tests a model’s ability to apply learned concepts and reasoning patterns to unseen scenarios.
-In our controlled experiments, both transformation and elements could be novel.
-Following this, we decompose task generalization into two primary dimensions: element generalization and transformation generalization.
+For fine-tuning state-of-the-art LLMs, we use a per-device batch size of 16 with 8 gradient accumulation steps (effective batch size of 128) and train for 24K optimization steps.
+We set the learning rate to 1e-4 and adopt a cosine learning-rate schedule with a 10% warm-up ratio.
+All fine-tuning experiments are conducted with bfloat16 (bf16) mixed-precision training.
+Given the scale of the data, the results of controlled experiments are averaged over three independent runs.
+Table 3: Hyperparameter settings for LLMs with different model sizes and architectures
+Architecture # Params Hidden Size Intermediate Size # Layers # Heads
+GPT 68K 32 N/A 4 4
+589K 80 N/A 7 8
+4.8M 256 N/A 6 4
+35M 512 N/A 11 8
+540M 1536 N/A 19 24
+3B 3072 N/A 26 32
+LLaMA 62K 48 128 2 4
+631K 80 216 8 4
+6M 288 768 6 6
+60M 640 1728 12 10
+623M 1536 4096 22 12
+3B 3072 8192 26 24
+D.2 Computational Cost
+We conduct training, fine-tuning, and inference on utilizing 8 NVIDIA A100 GPUs (80 GB memory) and 4 NVIDIA H200 GPUs.
+Appendix E Additional Theory and Propositions
+E.1 Task Distribution
+To investigate the extent to which CoT reasoning can handle tasks under various distribution discrepancies, we design task generalization experiments.
+As we discussed in Section 4, we decompose tasks into a combination of various transformations and elements.
+Therefore, we consider task generalization from two dimensions: transformation generalization and element generalization.
 Task Generalization Complexity. Guided by the data distribution lens, we first introduce a measure for generalization difficulty:
-Proposition 5.1 (Task Generalization Complexity).
+Proposition E.1 (Task Generalization Complexity).
 For a reasoning chain $f_{S}$ operating on elements $\mathbf{e}=(a_{0},\ldots,a_{l-1})$, define:
-$\operatorname{TGC}(C)=$ $\alpha\sum_{i=1}^{m}\mathbb{I}\left[a_{i}\notin\mathcal{E}^{i}_{\text{train }}\right]+\beta\sum_{j=1}^{n}\mathbb{I}\left[f_{j}\notin\mathcal{F}_{\text{train }}\right]+\gamma\mathbb{I}\left[\left(f_{1},f_{2},\ldots,f_{k}\right)\notin\mathcal{P}_{\text{train }}\right]+C_{T}$ (11)
+$\operatorname{TGC}(C)=\ \alpha\sum_{i=1}^{m}\mathbb{1}\left[a_{i}\notin\mathcal{E}^{i}_{\text{train }}\right]+\beta\sum_{j=1}^{n}\mathbb{1}\left[f_{j}\notin\mathcal{F}_{\text{train }}\right]+\gamma\mathbb{1}\left[\left(f_{1},f_{2},\ldots,f_{k}\right)\notin\mathcal{P}_{\text{train }}\right]+C_{T}$ (14)
 […]
 We establish a critical threshold beyond which CoT reasoning fails exponentially:
-Theorem 5.1 (Task Generalization Failure Threshold).
+Theorem E.1 (Task Generalization Failure Threshold).
 There exists a threshold $\tau$ such that when $\text{TGC}(C)>\tau$, the probability of correct CoT reasoning drops exponentially:
-$P(\text{correct}|C)\leq e^{-\delta(\text{TGC}(C)-\tau)}$ (12)
-The proof is provided in Appendix A.2.
-5.1 Transformation Generalization
-Transformation generalization evaluates the ability of CoT reasoning to effectively transfer when models encounter novel transformations during testing, which is an especially prevalent scenario in real-world applications.
-Figure 3: Performance of CoT reasoning on transformation generalization. Efficacy of CoT reasoning declines as the degree of distributional discrepancy increases.
-Experimental Setup.
-To systematically evaluate the impact of transformations, we conduct experiments by varying transformations between training and testing sets while keeping other factors constant (e.g., elements, length, and format).
-Guided by the intuition formalized in Proposition 5.1, we define four incremental levels of distribution shift in transformations as shown in Figure 2: (i) In-Distribution (ID): The transformations in the test set are identical to those observed during training, e.g., $f_{1}\circ f_{1}\rightarrow f_{1}\circ f_{1}$.
-(ii) Composition (CMP): Test samples comprise novel compositions of previously encountered transformations, though each individual transformation remains familiar, e.g., ${f_{1}\circ f_{1},f_{1}\circ f_{2},f_{2}\circ f_{1}}\rightarrow f_{2}\circ f_{2}$.
-(iii) Partial Out-of-Distribution (POOD): Test data include compositions involving at least one novel transformation not seen during training, e.g., $f_{1}\circ f_{1}\rightarrow f_{1}\circ f_{2}$.
-(iv) Out-of-Distribution (OOD): The test set contains entirely novel transformation types that are unseen in training, e.g., $f_{1}\circ f_{1}\rightarrow f_{2}\circ f_{2}$.
-Table 1: Full chain evaluation under different scenarios for transformation generalization.
-Transformation (Train $\rightarrow$ Test) Scenario Exact Match Edit Distance BLEU Score
-$f_{1}\circ f_{1}\rightarrow f_{1}\circ f_{1}$ ID 100.00% 0 1
-$\{f_{2}\circ f_{2},f_{1}\circ f_{2},f_{2}\circ f_{1}\}\rightarrow f_{1}\circ f_{1}$ CMP 0.01% 0.1326 0.6867
-$f_{1}\circ f_{2}\rightarrow f_{1}\circ f_{1}$ POOD 0.00% 0.1671 0.4538
-$f_{2}\circ f_{2}\rightarrow f_{1}\circ f_{1}$ OOD 0.00% 0.2997 0.2947
-Findings.
-Figure 3 illustrates the performance of the full chain under different distribution discrepancies computed by task generalize complexities (normalized between 0 and 1) in Definition 5.1.
-We can observe that, in general, the effectiveness of CoT reasoning decreases when distribution discrepancy increases.
+$P(C)\leq e^{-\delta(\text{TGC}(C)-\tau)}$ (15)
+The proof is provided in Appendix H.2.
+E.2 Length Distribution
+Length generalization examines how CoT reasoning degrades when models encounter test cases that differ in length from their training distribution.
+The difference in length could be introduced from the text space or the reasoning space of the problem.
+Therefore, we decompose length generalization into two complementary aspects: text length generalization and reasoning step generalization.
+Guided by instinct, we first propose to measure the length discrepancy.
+Proposition E.2 (Length Extrapolation Gaussian Degradation).
+For a model trained on chain-of-thought sequences of fixed length $L_{\text{train}}$, the generalization error at test length $L$ follows a Gaussian distribution:
+$\mathcal{E}(L)=\mathcal{E}_{0}+\left(1-\mathcal{E}_{0}\right)\cdot\left(1-\exp\left(-\frac{\left(L-L_{\text{train }}\right)^{2}}{2\sigma^{2}}\right)\right)$ (16)
+where $\mathcal{E}_{0}$ is the in-distribution error at $L=L_{\text{train}}$, $\sigma$ is the length generalization width parameter, and $L$ is the test sequence length
+The proof is provided in Appendix H.3.
+E.3 Format Distribution
+Format generalization assesses the robustness of CoT reasoning to surface-level variations in test queries.
+This dimension is especially crucial for determining whether models have internalized flexible, transferable reasoning strategies or remain reliant on the specific templates and phrasings encountered during training.
+Format Alignment Score. We introduce a metric for measuring prompt similarity:
+Definition E.1 (Format Alignment Score).
+For training prompt distribution $P_{train}$ and test prompt $p_{test}$:
+$\text{PAS}(p_{test})=\max_{p\in P_{train}}\cos(\phi(p),\phi(p_{test}))$ (17)
+where $\phi$ is a prompt embedding function.
+Table 4: Evaluation on transformation generalization.
+Transformation (Train $\rightarrow$ Test) Exact Match (%) Edit Distance BLEU Score
+Reasoning Answer Full Chain Full Chain Reasoning Answer Full Chain Reasoning Answer
+$\{f_{2}\circ f_{3},\;f_{3}\circ f_{2},\;f_{3}\circ f_{3}\}\rightarrow f_{2}\circ f_{2}$ 6.66 10.25 6.66 0.0941 0.0718 0.2244 0.5417 0.6683 0.1982
+$\{f_{2}\circ f_{3},\;f_{3}\circ f_{2},\;f_{2}\circ f_{2}\}\rightarrow f_{3}\circ f_{3}$ 100.00 9.19 9.19 0.0488 0.0000 0.1768 0.8220 1.0000 0.1932
+$f_{2}\circ f_{3}\rightarrow f_{3}\circ f_{2}$ 0.00 0.00 0.00 0.2997 0.3728 0.4808 0.2000 0.0019 0.0000
+$f_{3}\circ f_{2}\rightarrow f_{2}\circ f_{3}$ 0.00 0.00 0.00 0.2334 0.2249 0.4808 0.2548 0.0952 0.0000
+Appendix F Additional Quantitative Results
+F.1 Transformation Generalization
+F.1.1 Detailed analysis.
 For the instance shown in Table 1, from in-distribution to composition, POOD, and OOD, the exact match decreases from 1 to 0.01, 0, and 0, and the edit distance increases from 0 to 0.13, 0.17 when tested on data with transformation $f_{1}\circ f_{1}$.
@@
 As shown in Table 2, from $f_{1}\circ f_{2}$ to $f_{2}\circ f_{2}$, the LLMs can correctly answer 0.1% of questions.
-A close examination reveals that it is a coincidence, e.g., the query element is A, N, A, N, which happened to produce the same result for the two operations detailed in the Appendix D.1.
+A close examination reveals that it is a coincidence, e.g., the query element is A, N, A, N, which happened to produce the same result for the two operations detailed in the Appendix G.
 When further analysis is performed by breaking the full chain into reasoning steps and answers, we observe strong consistency between the reasoning steps and answers.
 For example, under the composition generalization setting, the reasoning steps are entirely correct on test data distribution $f_{1}\circ f_{1}$ and $f_{2}\circ f_{2}$, but with wrong answers.
-Probe these insistent cases in Appendix D.1, we can find that when a novel transformation (say $f_{1}\circ f_{1}$) is present, LLMs try to generalize the reasoning paths based on the most similar ones (i.e., $f_{1}\circ f_{2}$) seen during training, which leads to correct reasoning paths, yet incorrect answer, which echo the example in the introduction.
+Probe these insistent cases in Appendix G, we can find that when a novel transformation (say $f_{1}\circ f_{1}$) is present, LLMs try to generalize the reasoning paths based on the most similar ones (i.e., $f_{1}\circ f_{2}$) seen during training, which leads to correct reasoning paths, yet incorrect answer, which echo the example in the introduction.
 Similarly, generalization from $f_{1}\circ f_{2}$ to $f_{2}\circ f_{1}$ or vice versa allows LLMs to produce correct answers that are attributed to the commutative property between the two orthogonal transformations with unfaithful reasoning paths.
@@
 Rather than demonstrating a true understanding of text, CoT reasoning under task transformations appears to reflect a replication of patterns learned during training.
-Table 2: Evaluation on different components in CoT reasoning on transformation generalization. CoT reasoning shows inconsistency with the reasoning steps and answers.
-Transformation (Train $\rightarrow$ Test) Exact Match Edit Distance
-Reason Answer Full Chain Reason Answer Full Chain
-$\{f_{1}\circ f_{1},f_{1}\circ f_{2},f_{2}\circ f_{1}\}\rightarrow f_{2}\circ f_{2}$ 100.00% 0.01% 0.01% 0.000 0.481 0.133
-$\{f_{1}\circ f_{2},f_{2}\circ f_{1},f_{2}\circ f_{2}\}\rightarrow f_{1}\circ f_{1}$ 100.00% 0.01% 0.01% 0.000 0.481 0.133
-$f_{1}\circ f_{2}\rightarrow f_{2}\circ f_{1}$ 0.00% 100.00% 0.00% 0.373 0.000 0.167
-$f_{2}\circ f_{1}\rightarrow f_{1}\circ f_{2}$ 0.00% 100.00% 0.00% 0.373 0.000 0.167
-Figure 4: Performance on unseen transformation using SFT in various levels of distribution shift.
-Introducing a small amount of unseen data helps CoT reasoning to generalize across different scenarios.
-Experiment settings.
-To further probe when CoT reasoning can generalize to unseen transformations, we conduct supervised fine-tuning (SFT) on a small portion $\lambda$ of unseen data.
-In this way, we can decrease the distribution discrepancy between the training and test sets, which might help LLMs to generalize to test queries.
-Findings.
-As shown in Figure 4, we can find that generally a very small portion ($\lambda=1.5e^{-4}$) of data can make the model quickly generalize to unseen transformations.
-The less discrepancy between the training and testing data, the quicker the model can generalize.
-This indicates that a similar pattern appears in the training data, helping LLMs to generalize to the test dataset.
-5.2 Element Generalization
+F.1.2 Introducing another transformation.
+Our main conclusion is that CoT reasoning cannot generalize to genuinely novel transformations, including unseen compositions, even when the underlying primitives are well learned.
+Counterexamples based on commutativity have already substantiated this claim (Table 2).
+We now extend the analysis by introducing a non-commutative transformation $f_{3}$ and evaluating generalization behaviors that cannot be trivially explained by commutative equivalence.
+The results are summarized in Table 4.
+When models are trained on mixtures of transformations involving $f_{2}$ and $f_{3}$ and evaluated on unseen compositions (e.g., ${f_{2}\circ f_{3},f_{3}\circ f_{2},f_{3}\circ f_{3}}\rightarrow f_{2}\circ f_{2}$), performance remains extremely poor across all metrics.
+Exact match accuracy for the full chain stays below 10%, and both edit distance and BLEU scores indicate substantial divergence from the correct reasoning traces and final answers.
+Notably, even when the reasoning component occasionally achieves high exact match (e.g., 100% reasoning accuracy in ${f_{2}\circ f_{3},f_{3}\circ f_{2},f_{2}\circ f_{2}}\rightarrow f_{3}\circ f_{3}$), the corresponding full-chain and answer-level accuracy collapse, revealing a clear disconnect between locally plausible reasoning steps and globally correct execution.
+More strikingly, in strictly non-commutative transfer settings such as $f_{2}\circ f_{3}\rightarrow f_{3}\circ f_{2}$ and its reverse, the model fails completely: exact match drops to 0 across reasoning, answer, and full chain, while edit distance sharply increases and BLEU scores approach zero.
+Unlike earlier commutativity-induced cases—where incorrect reasoning paths could still yield correct answers—these failures demonstrate that once superficial equivalences are removed, CoT reasoning no longer exhibits any meaningful transfer.
+This provides strong evidence that prior apparent generalization was not driven by learning transformation semantics, but rather by exploiting distributional artifacts such as commutativity.
+Overall, Table 4 reinforces our central claim: CoT reasoning does not support systematic generalization to novel transformations.
+Instead, its success hinges on structural overlap and distributional shortcuts present in the training data.
+When these shortcuts are eliminated via non-commutative transformations, both reasoning traces and answers degrade simultaneously, exposing the brittleness of CoT reasoning under genuine task-level distribution shifts.
+Illustrative examples of $f_{3}$ are provided in Appendix B for completeness
+F.2 Element Generalization
+Figure 11: Element generalization results on various scenarios and relations.
 Element generalization is another critical factor to consider when LLMs try to generalize to new tasks.
@@
 While for OOD, atoms that constitute the elements are totally unseen during the training.
-Figure 5: Element generalization results on various scenarios and relations.
+(a) Performance on unseen element via SFT in various CMP scenarios.
+(b) Evaluation of CoT reasoning in SFT.
+Figure 12: SFT performances for element generalization. SFT helps to generalize to novel elements.
 Findings.
@@
 Most strikingly, the BLEU score is 0 when transferred to $f_{1}$ and $f_{2}$ transformations.
-A failure case in Appendix D.1 shows that the models cannot respond to any words when novel elements are present.
+A failure case in Appendix G shows that the models cannot respond to any words when novel elements are present.
 We further explore when CoT reasoning can generalize to novel elements by conducting SFT.
@@
 We find that there is a mismatch of accuracy between the answer and the reasoning step during the training process, which somehow might provide an explanation regarding why CoT reasoning is inconsistent in some cases.
-(a) Performance on unseen element via SFT in various CMP scenarios.
-(b) Evaluation of CoT reasoning in SFT.
-Figure 6: SFT performances for element generalization. SFT helps to generalize to novel elements.
-6 Length Generalization
-Length generalization examines how CoT reasoning degrades when models encounter test cases that differ in length from their training distribution.
-The difference in length could be introduced from the text space or the reasoning space of the problem.
-Therefore, we decompose length generalization into two complementary aspects: text length generalization and reasoning step generalization.
-Guided by instinct, we first propose to measure the length discrepancy.
-Length Extrapolation Bound. We establish a power-law relationship for length extrapolation:
-Proposition 6.1 (Length Extrapolation Gaussian Degradation).
-For a model trained on chain-of-thought sequences of fixed length $L_{\text{train}}$, the generalization error at test length $L$ follows a Gaussian distribution:
-$\mathcal{E}(L)=\mathcal{E}_{0}+\left(1-\mathcal{E}_{0}\right)\cdot\left(1-\exp\left(-\frac{\left(L-L_{\text{train }}\right)^{2}}{2\sigma^{2}}\right)\right)$ (13)
-where $\mathcal{E}_{0}$ is the in-distribution error at $L=L_{\text{train}}$, $\sigma$ is the length generalization width parameter, and $L$ is the test sequence length
-The proof is provided in Appendix A.3.
-Table 3: Evaluation for text length generalization.
-Length Exact Match (%) Edit Distance BLEU Score
-Full Chain Reason Answer Full Chain Reason Answer Full Chain Reason Answer
-2 0.00% 0.00% 0.00% 0.3772 0.4969 0.5000 0.4214 0.1186 0.0000
-3 0.00% 0.00% 0.00% 0.2221 0.3203 0.2540 0.5471 0.1519 0.0000
-4 100.00% 100.00% 100.00% 0.0000 0.0000 0.0000 1.0000 1.0000 1.0000
-5 0.00% 0.00% 0.00% 0.1818 0.2667 0.2000 0.6220 0.1958 0.2688
-6 0.00% 0.00% 0.00% 0.3294 0.4816 0.3337 0.4763 0.1174 0.2077
-6.1 Text Length Generalization
+Table 5: Evaluation on text length generalization.
+Text Length Exact Match (%) Edit Distance BLEU Score
+Reasoning Answer Full Chain Full Chain Reasoning Answer Full Chain Reasoning Answer
+2 0.00 0.00 0.00 0.3772 0.4969 0.5000 0.4214 0.1186 0.0000
+3 0.00 0.00 0.00 0.2221 0.3203 0.2540 0.5471 0.1519 0.0000
+4 100.00 100.00 100.00 0.0000 0.0000 0.0000 1.0000 1.0000 1.0000
+5 0.00 0.00 0.00 0.1818 0.2667 0.2000 0.6220 0.1958 0.2688
+6 0.00 0.00 0.00 0.3294 0.4816 0.3337 0.4763 0.1174 0.2077
+F.3 Text Length Generalization
 Text length generalization evaluates how CoT performance varies when the input text length (i.e., the element length $l$) differs from training examples.
@@
 For instance, from data with $l=4$ to those with $l=3$ or $l=5$, the BLEU score decreases from 1 to 0.55 and 0.62.
-Examples in Appendix D.1 indicate that LLMs attempt to produce CoT reasoning with the same length as the training data by adding or removing tokens in the reasoning chains.
+Examples in Appendix G indicate that LLMs attempt to produce CoT reasoning with the same length as the training data by adding or removing tokens in the reasoning chains.
 The efficacy of CoT reasoning length generalization deteriorates as the discrepancy increases.
@@
 However, the performance increases when we replace the padding with text by using the group strategy, which indicates its effectiveness.
-6.2 Reasoning Step Generalization
-The reasoning step generalization investigates whether models can extrapolate to reasoning chains requiring different steps $k$ from those observed during training.
-which is a popular setting in multi-step reasoning tasks.
-Experiment settings. Similar to text length generalization, we first pre-train the LLM with reasoning step $k=2$, and evaluate on data with reasoning step $k=1$ or $k=3$.
-(a) Reasoning step. From k=2 to k=1
-(b) Reasoning step. From k=2 to k=3
-Figure 8: Test performance for reasoning-step generalization across varying training data compositions. Performance varies systematically with changes in the distribution of training data components.
-Findings.
-As showcased in Figure 8, CoT reasoning cannot generalize across data requiring different reasoning steps, indicating the failure of generalization.
-Then, we try to decrease the distribution discrepancy introduced by gradually increasing the ratio of unseen data while keeping the dataset size the same when pre-training the model.
-And then, we evaluate the performance on two datasets.
-As we can observe, the performance on the target dataset increases along with the ratio.
-At the same time, the LLMs can not generalize to the original training dataset because of the small amount of training data.
-The trend is similar when testing different-step generalization, which follows the intuition and validates our hypothesis directly.
-7 Format Generalization
-Format generalization assesses the robustness of CoT reasoning to surface-level variations in test queries.
-This dimension is especially crucial for determining whether models have internalized flexible, transferable reasoning strategies or remain reliant on the specific templates and phrasings encountered during training.
-Format Alignment Score. We introduce a metric for measuring prompt similarity:
-Definition 7.1 (Format Alignment Score).
-For training prompt distribution $P_{train}$ and test prompt $p_{test}$:
-$\text{PAS}(p_{test})=\max_{p\in P_{train}}\cos(\phi(p),\phi(p_{test}))$ (14)
-where $\phi$ is a prompt embedding function.
-(a) Format generalization. Performance under various perturbation methods.
-(b) Format generalization. Performance vs. various applied perturbation areas.
-Figure 9: Performance of format generalization. Testing performance varies with different noise levels and areas where the noise is applied.
-Experiment settings.
-To systematically probe this, we introduce four distinct perturbation modes to simulate scenario in real-world: (i) insertion, where a noise token is inserted before each original token; (ii) deletion: it deletes the original token; (iii) modification: it replaces the original token with a noise token; and (iv) hybrid mode: it combines multiple perturbations.
-Each mode is applied for tokens with probabilities $p$, enabling us to quantify the model’s resilience to increasing degrees of prompt distribution shift.
-Findings.
-As shown in Figure 9a, we found that generally CoT reasoning can be easily affected by the format changes.
-No matter insertion, deletion, modifications, or hybrid mode, it creates a format discrepancy that affects the correctness.
-Among them, the deletion slightly affects the performance.
-While the insertions are relatively highly influential on the results.
-We further divide the query into several sections: elements, transformations, and prompt tokens.
-As shown in Figure 9b, we found that the elements and transformation play an important role in the format, whereas the changes to other tokens rarely affect the results.
-8 Temperature and Model Size
+F.4 Temperature and Model Size
 Temperature and model size generalization explores how variations in sampling temperature and model capacity can influence the stability and robustness of CoT reasoning.
@@
 Figure 14: Temperature and model size. The findings hold under different temperatures and model sizes.
+Figure 15: Task, length, and format generalization of LLMs with settings. The data distribution lens is invariant across LLMs with various sizes and architectures.
+Table 6: Task generalization performance of SOTA LLMs (mean $\pm$ std).
+Model Scenario Exact Match (%) Edit Distance BLEU Score
+LLaMA3-8B ID $100.00\pm 0.00$ $0.00\pm 0.00$ $1.00\pm 0.00$
+CMP $8.52\pm 0.00$ $0.23\pm 0.00$ $0.61\pm 0.00$
+POOD $0.00\pm 0.01$ $0.25\pm 0.01$ $0.46\pm 0.00$
+OOD $0.00\pm 0.00$ $0.27\pm 0.01$ $0.27\pm 0.00$
+Qwen3-14B-Instruct ID $100.00\pm 0.00$ $0.00\pm 0.00$ $1.00\pm 0.00$
+CMP $0.01\pm 0.01$ $0.17\pm 0.02$ $0.61\pm 0.00$
+POOD $0.00\pm 0.00$ $0.26\pm 0.01$ $0.42\pm 0.00$
+OOD $0.00\pm 0.00$ $0.38\pm 0.01$ $0.36\pm 0.00$
 Experiment settings.
@@
 Fig.
-10b shows the accuracy of models with different sizes using different SFT ratios, which closely matches the result of our default model size across all evaluated settings and configurations.
-9 Discussion and Implication
-Our investigation, conducted through the controlled environment of DataAlchemy, reveals that the apparent reasoning prowess of Chain-of-Thought (CoT) is largely a brittle mirage.
-The findings across task, length, and format generalization experiments converge on a conclusion: CoT is not a mechanism for genuine logical inference but rather a sophisticated form of structured pattern matching, fundamentally bounded by the data distribution seen during training.
-When pushed even slightly beyond this distribution, its performance degrades significantly, exposing the superficial nature of the “reasoning” it produces.
-While our experiments utilized models trained from scratch in a controlled environment, the principles uncovered are extensible to large-scale pre-trained models.
-We summarize the implications for practitioners as follows.
-Guard Against Over-reliance and False Confidence.
-CoT should not be treated as a “plug-and-play” module for robust reasoning, especially in high-stakes domains like medicine, finance, or legal analysis.
-The ability of LLMs to produce “fluent nonsense”—plausible but logically flawed reasoning chains—can be more deceptive and damaging than an outright incorrect answer, as it projects a false aura of dependability.
-Sufficient auditing from domain experts is indispensable.
-Prioritize Out-of-Distribution (OOD) Testing.
-Standard validation practices, where the test set closely mirrors the training set, are insufficient to gauge the true robustness of a CoT-enabled system.
-Practitioners must implement rigorous adversarial and OOD testing that systematically probes for vulnerabilities across task, length, and format variations.
-Recognize Fine-Tuning as a Patch, Not a Panacea.
-Our results show that Supervised Fine-Tuning (SFT) can quickly “patch” a model’s performance on a new, specific data distribution.
-However, this should not be mistaken for achieving true generalization.
-It simply expands the model’s “in-distribution” bubble slightly.
-Relying on SFT to fix every OOD failure is an unsustainable and reactive strategy that fails to address the core issue: the model’s lack of abstract reasoning capability.
-10 Conclusion
-In this paper, we critically examine the COT reasoning of LLMs through the lens of data distribution, revealing that the perceived structured reasoning capability largely arises from inductive biases shaped by in-distribution training data.
-We propose a controlled environment, DataAlchemy, allowing systematic probing of CoT reasoning along three crucial dimensions: task structure, reasoning length, and query format.
-Empirical findings consistently demonstrate that CoT reasoning effectively reproduces reasoning patterns closely aligned with training distributions but suffers significant degradation when faced with distributional deviations.
-Such observations reveal the inherent brittleness and superficiality of current CoT reasoning capabilities.
-We provide insights that emphasize real-world implications for both practitioners and researchers.
-[…]
-Appendix A Proof of Theorems
-A.1 Proof of CoT Generalization Bound
+14b shows the accuracy of models with different sizes using different SFT ratios, which closely matches the result of our default model size across all evaluated settings and configurations.
+Table 7: Reasoning step generalization of SOTA LLMs (mean $\pm$ std).
+Model Reasoning Step ($k$) Exact Match (%) Edit Distance BLEU Score
+LLaMA3-8B 1 $0.00\pm 0.00$ $0.75\pm 0.01$ $0.18\pm 0.00$
+2 $100.00\pm 0.00$ $0.00\pm 0.00$ $1.00\pm 0.00$
+3 $0.00\pm 0.00$ $0.54\pm 0.01$ $0.40\pm 0.00$
+Qwen3-14B-Instruct 1 $0.00\pm 0.00$ $0.54\pm 0.02$ $0.35\pm 0.00$
+2 $100.00\pm 0.00$ $0.00\pm 0.00$ $1.00\pm 0.00$
+3 $0.26\pm 0.05$ $0.65\pm 0.08$ $0.20\pm 0.00$
+Table 8: Format generalization under different noise levels for SOTA LLMs (mean $\pm$ std).
+Model Noise Level (%) Exact Match (%) Edit Distance BLEU Score
+LLaMA3-8B 0 $100.00\pm 0.00$ $0.00\pm 0.00$ $1.00\pm 0.00$
+5 $32.30\pm 0.47$ $41\pm 0.02 0.0$$04\pm 0.00$
+10 $17.74\pm 0.38$ $45\pm 0.00 0.0$$04\pm 0.00$
+15 $9.27\pm 0.29$ $49\pm 0.05 0.0$$04\pm 0.00$
+20 $4.72\pm 0.21$ $53\pm 0.03 0.0$$03\pm 0.00$
+25 $2.32\pm 0.15$ $57\pm 0.00 0.0$$03\pm 0.00$
+30 $1.09\pm 0.10$ $60\pm 0.08 0.0$$03\pm 0.00$
+Qwen3-14B-Instruct 0 $100.00\pm 0.00$ $0.00\pm 0.00$ $1.00\pm 0.00$
+5 $9.96\pm 0.30$ $0.22\pm 0.03$ $0.41\pm 0.00$
+10 $5.09\pm 0.22$ $0.35\pm 0.01$ $0.24\pm 0.00$
+15 $2.43\pm 0.15$ $0.44\pm 0.02$ $0.16\pm 0.00$
+20 $1.19\pm 0.11$ $0.52\pm 0.05$ $0.11\pm 0.00$
+25 $0.53\pm 0.07$ $57\pm 0.04 0.0$$08\pm 0.00$
+30 $0.22\pm 0.05$ $62\pm 0.02 0.0$$06\pm 0.00$
+F.5 Internal Validity
+Figures 8 and 15 illustrate task, length, and format generalization across a wide range of GPT- and LLaMA-style models.
+Increasing distribution discrepancy leads to a monotonic degradation of CoT performance regardless of architectural choice or parameter count.
+While larger models achieve uniformly higher absolute scores under near in-distribution conditions, they do not exhibit qualitatively different robustness profiles under moderate or severe shifts.
+This suggests that the observed failures of CoT reasoning cannot be attributed to insufficient capacity or architectural idiosyncrasies, but rather reflect a shared inductive bias learned from training distributions.
+Model scaling consistently improves performance in ID and mildly shifted regimes (e.g., CMP or small reasoning-step extrapolation), but provides diminishing returns as the discrepancy increases toward POOD and OOD settings.
+In particular, larger models tend to preserve fluent intermediate reasoning traces even when final answers deteriorate, mirroring the same failure modes observed in smaller models.
+This pattern reinforces the interpretation that scaling primarily enhances pattern interpolation within the support of the training distribution, rather than enabling principled extrapolation beyond it.
+Notably, the qualitative nature of errors remains stable across model sizes and architectures.
+Under task shifts, models consistently default to the closest seen transformation pattern; under length shifts, they bias toward producing training-length reasoning chains; and under format perturbations, they remain sensitive to surface-level noise in structurally salient regions.
+The persistence of these behaviors across settings further supports the internal validity of the data distribution lens: CoT reasoning behaves as a distribution-sensitive generative process rather than an architecture-specific reasoning mechanism.
+Taken together, these results demonstrate that our core findings are not artifacts of a particular model family, scale, or training instability.
+Instead, the dependence of CoT effectiveness on task, length, and format distributions emerges as a stable and reproducible phenomenon across controlled LLM instantiations.
+This strengthens the claim that distribution discrepancy—rather than model choice—is the dominant factor governing when CoT reasoning succeeds or fails.
+F.6 External Validity
+As illustrated in Table 6, 7 & 8, the results on LLaMA3-8B-Instruct and Qwen3-14B-Instruct confirm that the behaviors identified in DataAlchemy persist in real-world, pretrained LLMs despite unknown and opaque training distributions.
+In the in-distribution (ID) setting, both models achieve perfect performance (100% exact match, zero edit distance, BLEU score of 1), indicating that the curated tasks are well within the expressive and optimization capacity of modern instruction-tuned models.
+However, once distribution shifts are introduced, performance degrades sharply and systematically.
+Under composition (CMP), exact match accuracy drops to 8.52% for LLaMA3-8B and near zero (0.01%) for Qwen3-14B, while POOD and OOD settings result in complete failure (0% exact match) for both models.
+Correspondingly, edit distance increases and BLEU score decreases monotonically from CMP to OOD.
+These trends mirror those observed in models trained from scratch, reinforcing that the failure of CoT reasoning under task-level distribution shifts is not an artifact of synthetic training or limited model scale.
+Reasoning-step generalization further highlights the limited extrapolation ability of CoT reasoning in SOTA models.
+Both LLaMA3-8B and Qwen3-14B achieve perfect performance at the in-distribution reasoning depth ($k=2$), but fail almost entirely at unseen depths.
+For $k=1$ and $k=3$, exact match accuracy collapses to 0% in nearly all cases, accompanied by large edit distances and sharply reduced BLEU scores.
+Notably, Qwen3-14B exhibits a marginal non-zero accuracy (0.26%) at $k=3$, but this gain is unstable and negligible relative to the ID performance.
+This pattern indicates that even large, instruction-tuned models do not acquire a length-agnostic or algorithmic reasoning procedure, but instead internalize a narrowly scoped reasoning template tied to the training distribution.
+Format generalization experiments show consistent degradation trends across noise levels for both models.
+As noise increases from 0% to 30%, exact match accuracy decays smoothly from 100% to near zero, while edit distance increases and BLEU score decreases monotonically.
+Although LLaMA3-8B initially appears more sensitive at low noise levels (e.g., 32.30% exact match at 5% noise versus 9.96% for Qwen3-14B), both models converge to similarly poor performance under higher noise.
+The near-parallel degradation curves suggest that robustness to surface-level perturbations is constrained by distributional alignment.
+Across all settings, the reported standard deviations are small relative to the absolute performance gaps between ID and shifted distributions.
+This indicates that the observed trends are stable across runs and not driven by sampling noise or stochastic decoding effects.
+In particular, zero or near-zero variance in ID settings confirms deterministic mastery of in-distribution patterns, while low variance under OOD conditions reflects consistently poor generalization rather than brittle or erratic behavior.
+Overall, these results establish strong external validity for the proposed data distribution lens.
+Despite their scale, architectural sophistication, and instruction tuning, SOTA LLMs exhibit the same qualitative behaviors as controlled models trained in DataAlchemy: strong in-distribution performance, sharp degradation under task, length, and format shifts, and limited benefit from increased model capacity.
+This alignment between synthetic and real-world settings supports the claim that the effectiveness and brittleness of CoT reasoning are fundamentally governed by distributional alignment, rather than by pretraining scale or architectural advances alone.
+Appendix G Additional Qualitative Analysis
+G.1 Failures in Task Generalization
+G.1.1 Failures caused by Transformation Generalization
+The following case illustrates the failure caused by the transformation partial OOD setting discussed in Section C.1.1.
+Specifically, the model is trained exclusively on the compositional transformation $f_{1}\circ f_{2}$, but is evaluated on the unseen composition $f_{1}\circ f_{1}$.
+Query:
+H U S P [F1] [F1] <think>
+Response:
+U H F C [F2] <answer> H F C U
+Ground-Truth:
+U H F C [F1] <answer> H U S P
+The results demonstrate that the model persists in applying the training composition ($f_{1}\circ f_{2}$) rather than adapting to the new sequence ($f_{1}\circ f_{1}$), even though the constituent transformation $f_{1}$ was observed during training.
+The model incorrectly generates the token [F2] in the reasoning step, mirroring the fixed pattern learned from the training data.
+G.1.2 Failures caused by Element Generalization
+The following case illustrates the failure caused by the element-level OOD setting discussed in Section C.1.2.
+Specifically, the training data was restricted to visible elements A–M, while the model was evaluated on queries consisting exclusively of elements N–Z.
+Query:
+N N N Q [F2] [F1] <think>
+Response:
+A A C E [F1] <answer> N N P R
+Ground-Truth:
+N N Q N [F1] <answer> A A D A
+In the generated reasoning, the model incorrectly outputs AACE instead of the expected intermediate result NNQN.
+This error arises because all training queries consisted solely of elements A–M.
+Consequently, the model fails to generalize the transformation $f_{2}$ to the unseen elements; instead, it reverts to the training distribution, attempting to replicate the A–M patterns observed during training.
+G.2 Failures in Length Generalization
+G.2.1 Failures in Text Length Generalization
+The following failure case demonstrates the model’s inability to generalize to unseen text lengths.
+In this experiment, the model was trained on sequences of length four but evaluated on sequences of length five.
+Query:
+I G L L Q [F1] [F2] <think>
+Response:
+T Y Y [F2] <answer> T Y Y V
+Ground-Truth:
+V T Y Y D [F2] <answer> T Y Y D V
+Despite the increased length of the input query (five atoms), the model fails to adapt.
+As seen in the generated answer (T Y Y V), the model rigidly adheres to the length constraint observed during training, outputting a sequence of four atoms instead of the required five.
+G.2.2 Failures in Reasoning Step Generalization
+The following case shows that model trained under $f_{1}\circ f_{1}$ tried to reproduce the length in training data by adding tokens in the reason chain even prompted with seen transformation $f_{1}$
+Query:
+A A B D [F1] <answer>
+Response:
+N O A Z N N O Q [F1] <answer> A A B D
+Ground-Truth:
+N N O Q
+G.3 Failures in Format Generalization
+The following failure case demonstrates the model’s inability to generalize under format change. Specifically, here we use delete as the format-changing mechanism.
+Query:
+A A A T [F1] <answer>
+Response:
+N G N G Y
+Ground-Truth:
+N N N G
+Appendix H Proof of Theorems
+H.1 Proof of CoT Generalization Bound
 Proof.
-Let $f_{\theta}$ be a model trained on samples from the distribution $\mathcal{D}_{\text{train}}$ using a loss function $\ell(f_{\theta}(x),y)$ that is $\Lambda$-Lipschitz and bounded.
-The expected test risk is given by
-$R_{\text{test}}(f_{\theta})=\mathbb{E}_{(x,y)\sim\mathcal{D}_{\text{test}}}\left[\ell(f_{\theta}(x),y)\right].$ (15)
-We can decompose the test risk as
-$R_{\text{test}}(f_{\theta})=R_{\text{train}}(f_{\theta})+\left(R_{\text{test}}(f_{\theta})-R_{\text{train}}(f_{\theta})\right).$ (16)
-To bound the discrepancy between $R_{\text{test}}$ and $R_{\text{train}}$, we invoke a standard result from statistical learning theory.
-Given that $\ell$ is $\Lambda$-Lipschitz and the discrepancy measure $\Delta(\mathcal{D}_{\text{train}},\mathcal{D}_{\text{test}})$ is an integral probability metric (e.g., Wasserstein-1 distance), we have
-$\left|R_{\text{test}}(f_{\theta})-R_{\text{train}}(f_{\theta})\right|\leq\Lambda\cdot\Delta(\mathcal{D}_{\text{train}},\mathcal{D}_{\text{test}}).$ (17)
-Therefore, the test risk satisfies
-$R_{\text{test}}(f_{\theta})\leq R_{\text{train}}(f_{\theta})+\Lambda\cdot\Delta(\mathcal{D}_{\text{train}},\mathcal{D}_{\text{test}}).$ (18)
-We next account for the generalization gap between the empirical training risk $\hat{R}_{\text{train}}(f_{\theta})$ and the expected training risk $R_{\text{train}}(f_{\theta})$.
-By applying a concentration inequality (e.g., Hoeffding’s inequality), with probability at least $1-\delta$, we have
-$R_{\text{train}}(f_{\theta})\leq\hat{R}_{\text{train}}(f_{\theta})+\mathcal{O}\left(\sqrt{\frac{\log(1/\delta)}{n}}\right),$ (19)
-where $n$ is the number of training samples.
-Combining the above, we obtain that with high probability,
-$R_{\text{test}}(f_{\theta})\leq\hat{R}_{\text{train}}(f_{\theta})+\Lambda\cdot\Delta(\mathcal{D}_{\text{train}},\mathcal{D}_{\text{test}})+\mathcal{O}\left(\sqrt{\frac{\log(1/\delta)}{n}}\right).$ (20)
-This concludes the proof. ∎
-A.2 Proof of Task Generalization Failure Threshold
+Recall that the empirical and population training risks and the test risk are
+$\hat{R}_{\text{train}}(f_{\theta})$ $=\frac{1}{n}\sum_{i=1}^{n}\ell\bigl(f_{\theta}(x_{i}),y_{i}\bigr),$ (18)
+$R_{\text{train}}(f_{\theta})$ $=\mathbb{E}_{(x,y)\sim\mathcal{D}_{\text{train}}}\bigl[\ell(f_{\theta}(x),y)\bigr]$ (19)
+$R_{\text{test}}(f_{\theta})$ $=\mathbb{E}_{(x,y)\sim\mathcal{D}_{\text{test}}}\bigl[\ell(f_{\theta}(x),y)\bigr]$ (20)
+with $(x_{i},y_{i})\sim\mathcal{D}_{\text{train}}$ i.i.d., and the loss satisfies $0\leq\ell(f_{\theta}(x),y)\leq B$ for all $(x,y)$. We also defined
+$\Delta(\mathcal{D}_{\text{train}},\mathcal{D}_{\text{test}}):=\operatorname{TV}(\mathcal{D}_{\text{train}},\mathcal{D}_{\text{test}}).$ (21)
+We decompose
+$R_{\text{test}}(f_{\theta})-\hat{R}_{\text{train}}(f_{\theta})=\ \bigl(R_{\text{test}}(f_{\theta})-R_{\text{train}}(f_{\theta})\bigr)+\bigl(R_{\text{train}}(f_{\theta})-\hat{R}_{\text{train}}(f_{\theta})\bigr)$ (22)
+and bound the two terms separately.
+First, let $P=\mathcal{D}_{\text{train}}$, $Q=\mathcal{D}_{\text{test}}$ and $g(x,y)=\ell(f_{\theta}(x),y)$. Using the fact that for any measurable $g$ with $|g|\leq B$,
+$\bigl|\mathbb{E}_{P}[g]-\mathbb{E}_{Q}[g]\bigr|\leq 2B\,\operatorname{TV}(P,Q)$ (23)
+we obtain
+$\bigl|R_{\text{test}}(f_{\theta})-R_{\text{train}}(f_{\theta})\bigr|\leq 2B\,\Delta(\mathcal{D}_{\text{train}},\mathcal{D}_{\text{test}})$ (24)
+and in particular,
+$R_{\text{test}}(f_{\theta})\leq R_{\text{train}}(f_{\theta})+2B\,\Delta(\mathcal{D}_{\text{train}},\mathcal{D}_{\text{test}})$ (25)
+Second, define $Z_{i}:=\ell(f_{\theta}(x_{i}),y_{i})\in[0,B]$. Then
+$\hat{R}_{\text{train}}(f_{\theta})=\frac{1}{n}\sum_{i=1}^{n}Z_{i},\ R_{\text{train}}(f_{\theta})=\mathbb{E}[Z_{i}]$ (26)
+By Hoeffding’s inequality, for any $\varepsilon>0$,
+$\Pr\Bigl(\hat{R}_{\text{train}}(f_{\theta})-R_{\text{train}}(f_{\theta})\geq\varepsilon\Bigr)\leq\exp\!\left(-\frac{2n\varepsilon^{2}}{B^{2}}\right)$ (27)
+Setting the right-hand side to $\delta$ and solving for $\varepsilon$ yields $\varepsilon=B\sqrt{\frac{\log(1/\delta)}{2n}}$, so with probability at least $1-\delta$,
+$R_{\text{train}}(f_{\theta})\leq\hat{R}_{\text{train}}(f_{\theta})+B\sqrt{\frac{\log(1/\delta)}{2n}}$ (28)
+Combining Equation (24) and Equation (28), we conclude that, with probability at least $1-\delta$,
+$R_{\text{test}}(f_{\theta})$ $\leq R_{\text{train}}(f_{\theta})+2B\,\Delta(\mathcal{D}_{\text{train}},\mathcal{D}_{\text{test}})$ (29)
+$\leq\hat{R}_{\text{train}}(f_{\theta})+B\sqrt{\frac{\log(1/\delta)}{2n}}+2B\,\Delta(\mathcal{D}_{\text{train}},\mathcal{D}_{\text{test}})$
+which completes the proof. ∎
+H.2 Proof of Task Generalization Failure Threshold
+Proof.
 We establish the exponential decay bound through a probabilistic analysis of reasoning failure modes in the presence of task generalization complexity.
@@
 Here we make the assumption that the reasoning failures induced by novel arguments, functions, and patterns contribute independently to the overall failure probability and hence we model the success probability as a product of component-wise success rates:
-$P(\text{correct}|C)=P_{0}\prod_{i=1}^{m}\rho_{a}^{\mathbb{I}[A_{i}]}\prod_{j=1}^{n}\rho_{f}^{\mathbb{I}[F_{j}]}\rho_{p}^{\mathbb{I}[\mathcal{Q}]}\rho_{c}^{C_{T}}$
+$P(C)=P_{0}\prod_{i=1}^{m}\rho_{a}^{\mathbb{1}[A_{i}]}\prod_{j=1}^{n}\rho_{f}^{\mathbb{1}[F_{j}]}\rho_{p}^{\mathbb{1}[\mathcal{Q}]}\rho_{c}^{C_{T}}$ (30)
 where $P_{0}\in(0,1]$ represents the baseline success probability when all components are within the training distribution, and $\rho_{a},\rho_{f},\rho_{p},\rho_{c}\in(0,1)$ are the degradation factors associated with novel arguments, functions, patterns, and task-specific complexity, respectively.
-$\ln P(\text{correct}\mid C)\;=$ $\ln P_{0}+\sum_{i=1}^{m}\mathbb{I}[A_{i}]\,\ln\rho_{a}+\sum_{j=1}^{n}\mathbb{I}[F_{j}]\,\ln\rho_{f}+\mathbb{I}[\mathcal{Q}]\,\ln\rho_{p}+C_{T}\,\ln\rho_{c}$ (21)
+$\ln P(C)=\ \ln P_{0}+\sum_{i=1}^{m}\mathbb{1}[A_{i}]\,\ln\rho_{a}+\sum_{j=1}^{n}\mathbb{1}[F_{j}]\,\ln\rho_{f}+\mathbb{1}[\mathcal{Q}]\,\ln\rho_{p}+C_{T}\,\ln\rho_{c}$ (31)
 For notational convenience, we define the positive constants:
-$\xi_{a}:=-\ln\rho_{a}>0,\xi_{f}:=-\ln\rho_{f}>0,\xi_{p}:=-\ln\rho_{p}>0,\xi_{c}:=-\ln\rho_{c}>0$
+$\xi_{a}=-\ln\rho_{a}>0,\ \xi_{f}=-\ln\rho_{f}>0,$ (32)
+$\xi_{p}=-\ln\rho_{p}>0,\ \xi_{c}=-\ln\rho_{c}>0$
 hence we have:
-$\ln P(\text{correct}|C)=\ln P_{0}-\xi_{a}\sum_{i=1}^{m}\mathbb{I}[A_{i}]-\xi_{f}\sum_{j=1}^{n}\mathbb{I}[F_{j}]-\xi_{p}\mathbb{I}[\mathcal{Q}]-\xi_{c}C_{T}$ (22)
+$\ln P(C)=$ $\ln P_{0}-\xi_{a}\sum_{i=1}^{m}\mathbb{1}[A_{i}]-\xi_{f}\sum_{j=1}^{n}\mathbb{1}[F_{j}]$ (33)
+$-\xi_{p}\mathbb{1}[\mathcal{Q}]-\xi_{c}C_{T}$
 Lemma: Relationship to TGC. The expression in equation above can be bounded in terms of $\text{TGC}(C)$ as follows:
-$\ln P(\text{correct}|C)\leq\ln P_{0}-\delta\cdot\text{TGC}(C)$ (23)
+$\ln P(C)\leq\ln P_{0}-\delta\cdot\text{TGC}(C)$ (34)
 where $\delta=\min(\frac{\xi_{a}}{\alpha},\frac{\xi_{f}}{\beta},\frac{\xi_{p}}{\gamma},\xi_{c})>0$.
 Proof of Lemma: From the definition of $\text{TGC}(C)$ in Eq. (14), we have:
-$\text{TGC}(C)=\alpha\sum_{i=1}^{m}\mathbb{I}[A_{i}]+\beta\sum_{j=1}^{n}\mathbb{I}[F_{j}]+\gamma\mathbb{I}[\mathcal{Q}]+C_{T}$ (24)
+$\text{TGC}(C)=\alpha\sum_{i=1}^{m}\mathbb{1}[A_{i}]+\beta\sum_{j=1}^{n}\mathbb{1}[F_{j}]+\gamma\mathbb{1}[\mathcal{Q}]+C_{T}$ (35)
 By the definition of $\delta$, each term in Eq. (33) satisfies:
-$\xi_{a}\sum_{i=1}^{m}\mathbb{I}[A_{i}]\geq\delta\alpha\sum_{i=1}^{m}\mathbb{I}[A_{i}]$ (25)
-$\xi_{f}\sum_{j=1}^{n}\mathbb{I}[F_{j}]\geq\delta\beta\sum_{j=1}^{n}\mathbb{I}[F_{j}]$ (26)
-$\xi_{p}\mathbb{I}[\mathcal{Q}]\geq\delta\gamma\mathbb{I}[\mathcal{Q}]$ (27)
+$\xi_{a}\sum_{i=1}^{m}\mathbb{1}[A_{i}]\geq\delta\alpha\sum_{i=1}^{m}\mathbb{1}[A_{i}]$ (36)
+$\xi_{f}\sum_{j=1}^{n}\mathbb{1}[F_{j}]\geq\delta\beta\sum_{j=1}^{n}\mathbb{1}[F_{j}]$ (37)
+$\xi_{p}\mathbb{1}[\mathcal{Q}]\geq\delta\gamma\mathbb{1}[\mathcal{Q}]$ (38)
 $\xi_{c}C_{T}\geq\delta C_{T}$ (39)
@@
 We now define the threshold $\tau:=\frac{\ln P_{0}}{\delta}$. From Eq. (34), when $\text{TGC}(C)>\tau$, we have:
-$\ln P(\operatorname{correct}\mid C)$ $\leq\ln P_{0}-\delta\cdot\operatorname{TGC}(C)$ (29)
-$=\delta(\tau-\operatorname{TGC}(C))$ (30)
-$=-\delta(\operatorname{TGC}(C)-\tau)$ (31)
-Exponentiating both sides yields the desired bound: $P(\operatorname{correct}\mid C)\leq e^{-\delta(\mathrm{TGC}(C)-\tau)}$
-A.3 Proof of Length Extrapolation Bound
+$\ln P(C)$ $\leq\ln P_{0}-\delta\cdot\operatorname{TGC}(C)$ (40)
+$=\delta(\tau-\operatorname{TGC}(C))$
+$=-\delta(\operatorname{TGC}(C)-\tau)$
+Exponentiating both sides yields the desired bound: $P(C)\leq e^{-\delta(\mathrm{TGC}(C)-\tau)}$
+H.3 Proof of Length Extrapolation Bound
 Proof.
@@
 ∎
-Appendix B Experiment Details
-The model is trained using the AdamW optimiser in mixed precision (FP16).
-The default learning rate is $3\times 10^{-3}$, and the schedule follows a cosine decay with a 10% warm-up ratio.
-Training is conducted for 10 epochs, using a batch size of 1024.
-A weight decay of 0.01 is applied, and gradient norms are clipped at 1.0.
-Appendix C Illustration of Datasets
-Below are the examples of transformation $f_{1}$ and $f_{2}$:
-Transformation[F1]: A A F Q [F1] <answer> N N S D
-Transformation[F2]: A A L P [F2] <answer> A L P A
-aside from single transformation, we can composite transformations arbitrarily:
-Transformation[F1F2]: A C I A [F1] [F2] <think>
-N P V N [F2] <answer>
-P V N N
-Transformation[$F2$]: N O V S [F2] [F2] <think>
-O V S N [F2] <answer>
-V S N O
-we use exact match, edit distance and BELU score to measure the discrepancy between generated tokens and the labels.
-For more than one transformation examples, we can further measure the discrepancy for reasoning and answering seperately.
-Appendix D Additional Experimental Results
-D.1 Additional Qualitative Analysis
-D.1.1 Orthognal Transformation Caused Coincidence
-The following case shows that even if the transformation is different, the model that trained on transformation $f_{2}\circ f_{1}$ can still provide correct answer through incorrect reasoning:
-Prompt: A A A B [F1] [F2] <answer>’
-Generated:’B A A A [F1] <answer> O N N N
-Expected: O N N N
-D.1.2 Correct reasoning but failed in final answer
-The following case shows that the model pretrained on the union of three transformation $f_{1}\circ f_{2},f_{2}\circ f_{1},f_{2}\circ f_{2}$ and test on $f_{1}\circ f_{1}$
-Prompt: A A A D [R1] [R1] <think>’
-Generated: N N N Q [R1] <answer> N N Q N
-Expected: N N N Q [R1] <answer> A A A D
-D.1.3 Failure to generalize to novel element
-The following case shows that the model trained on element set $a_{i}\in[A,M]$ can not generalize to unseen elements such as N or O
-Prompt: N N N O [F1] [F1] <think>’
-Generated: R V Q S [F1] <answer> E I D F
-Expected: A A A B [F1] <answer> N N N O
-D.1.4 LLM reproduces CoT reasoning at seen lengths
-The following case shows that model trained under $f_{1}\circ f_{1}$ tried to reproduce the length in training data by adding tokens in the reason chain even prompted with seen transformation $f_{1}$
-Prompt: A A B D [f1] <answer>’
-Generated: N O A Z N N O Q [f1]
-<answer> A A B D
-Expected: N N O Q
+Appendix I Discussion and Implication
+Our investigation, conducted through the controlled environment of DataAlchemy, reveals that the apparent reasoning prowess of Chain-of-Thought (CoT) is largely a brittle mirage.
+The findings across task, length, and format generalization experiments converge on a conclusion: CoT is not a mechanism for genuine logical inference but rather a sophisticated form of structured pattern matching, fundamentally bounded by the data distribution seen during training.
+When pushed even slightly beyond this distribution, its performance degrades significantly, exposing the superficial nature of the “reasoning” it produces.
+While our experiments utilized models trained from scratch in a controlled environment, the principles uncovered are extensible to large-scale pre-trained models.
+We summarize the implications for practitioners as follows.
+Guard against over-reliance and false confidence.
+CoT should not be treated as a “plug-and-play” module for robust reasoning, especially in high-stakes domains like medicine, finance, or legal analysis.
+The ability of LLMs to produce “fluent nonsense”—plausible but logically flawed reasoning chains—can be more deceptive and damaging than an outright incorrect answer, as it projects a false aura of dependability.
+Sufficient auditing from domain experts is indispensable.
+Prioritize OOD testing.
+Standard validation practices, where the test set closely mirrors the training set, are insufficient to gauge the true robustness of a CoT-enabled system.
+Practitioners must implement rigorous adversarial and OOD testing that systematically probes for vulnerabilities across task, length, and format variations.
+Recognize fine-Tuning as a patch, not a Panacea.
+Our results show that Supervised Fine-Tuning (SFT) can quickly “patch” a model’s performance on a new, specific data distribution.
+However, this should not be mistaken for achieving true generalization.
+It simply expands the model’s “in-distribution” bubble slightly.
+Relying on SFT to fix every OOD failure is an unsustainable and reactive strategy that fails to address the core issue: the model’s lack of abstract reasoning capability.
+Appendix J Use of Generative AI
+To enhance clarity and readability, we utilized the GPT-5.2 model exclusively as a language polishing tool.
+Its role was confined to proofreading, grammatical correction, and stylistic refinement—functions analogous to those provided by traditional grammar checkers and dictionaries.
+This tool did not contribute to the generation of new scientific content or ideas, and its usage is consistent with standard practices for manuscript preparation.